### v1

# Subnormalisers of semisimple elements in finite groups of Lie type

Abstract.
We determine subnormalisers of semisimple elements of prime power order in finite quasi-simple groups of Lie type. For this, we determine the maximal overgroups of normalisers of Sylow tori. This is motivated by the recent character correspondence conjecture by Moretó and Rizo as well as by the question of existence of quasi-semiregular elements in finite permutation groups.
Key words and phrases:
Subnormalisers, Sylow $d$-tori, torus normalisers, groups of Lie type, sporadic groups
2020 Mathematics Subject Classification:
20D06, 20D20, 20D35, 20E25, 20G40
The author gratefully acknowledges financial support by the DFG— Project-ID 286237555 – TRR 195. He also thanks Mandi Schaeffer Fry and Juan Martínez Madrid for their comments on an earlier version.
1. Introduction
In this paper we continue our investigation in [16] in relation to the recent conjecture of Moretó and Rizo [22] on character correspondences for finite groups $G$. For a prime $p$, these should relate the irreducible characters of $G$ to those of subnormalisers of $p$-elements of $G$, which are defined as follows. For a subgroup $H$ of $G$ let
$S_{G}(H):=\{g\in G\mid H{\,\triangleleft\triangleleft\,}\langle g,H\rangle\},$
the set of elements $g\in G$ such that $H$ is subnormal in $\langle g,H\rangle$, and define
${\operatorname{Sub}}_{G}(H):=\big\langle S_{G}(H)\big\rangle$
to be the subnormaliser of $H$. If $H=\langle x\rangle$ is generated by a single element, we also write ${\operatorname{Sub}}_{G}(x)$ for ${\operatorname{Sub}}_{G}(\langle x\rangle)$. For an element $x\in G$ let ${\operatorname{Irr}}^{x}(G)$ denote the set of complex irreducible characters of $G$ that do not vanish at $x$. The following was put forward in [22]:
Conjecture 1 (Moretó–Rizo).
Let $G$ be a finite group and $p$ a prime. Then for any $p$-element $x\in G$ there exists a bijection $f_{x}:{\operatorname{Irr}}^{x}(G)\to{\operatorname{Irr}}^{x}({\operatorname{Sub}}_{G}(x))$ such that
  1. (1)
    $\chi(1)_{p}=f_{x}(\chi)(1)_{p}$, and
  2. (2)
    ${\mathbb{Q}}(\chi(x))={\mathbb{Q}}(f_{x}(\chi)(x))$.
In order to investigate this conjecture for non-abelian simple groups it seems useful to understand the structure of subnormalisers of $p$-elements. In our predecessor paper [16] we classified semisimple picky $p$-elements, that is, elements whose subnormaliser is a Sylow $p$-normaliser, of quasi-simple groups of Lie type except for $p\leq 3$, and obtained partial information on subnormalisers of unipotent elements. The picky semisimple 2- and 3-elements were then determined in [17]. Here, we continue the investigation of subnormalisers for semisimple $p$-elements in groups of Lie type. This naturally leads to the question of understanding overgroups of certain torus normalisers which might be of independent interest.
Our main result is the determination of maximal overgroups of normalisers of Sylow $d$-tori, for integers $d\geq 1$, which in turn yields the maximal overgroups of normalisers of abelian Sylow subgroup, and using this the description of subnormalisers of semisimple elements lying in these abelian Sylow subgroups. It turns out, a posteriori, that these subnormalisers are natural geometrically defined subgroups except for one single case in the exceptional group $G_{2}(4)$ where the sporadic simple group $J_{2}$ of Janko appears. Combining our analysis with results for symmetric and sporadic groups we can state:
Theorem 2.
Let $G$ be a finite quasi-simple group and $p$ a prime such that $G$ has abelian Sylow $p$-subgroups. Then the subnormalisers of all $p$-elements in $G$ are known.
The situation for non-abelian Sylow subgroups is considerably more difficult with further types of subnormalisers appearing, and not so tightly related to Sylow tori, so we will not discuss it here.
Our investigations are related to other current research work. Subnormalisers play a central role in the investigation of Giudici, Morgan and Praeger on finite permutation groups $G$ containing quasi-semiregular elements, since $x\in G$ of prime order is quasi-semiregular if and only if there is a point-stabiliser of $G$ containing ${\operatorname{Sub}}_{G}(x)$ (see [10, Thm 3.3]).
Baumeister, Burness, Guralnick and Tong-Viet [1] classify finite non-abelian almost simple groups with a Sylow $p$-subgroup contained in a unique maximal subgroup. This is related to our work for simple groups of Lie type $G$ with an abelian Sylow $p$-subgroup $P$ as follows. If $P$ is contained in a unique maximal subgroup $M$ of $G$, then so is its normaliser ${\mathbf{N}}_{G}(P)$; in particular if this normaliser lies in several maximal overgroups, the same is true a fortiori for $P$. That is, all examples in [1] with abelian Sylow subgroups also show up as part of our classification.
Our paper is built up as follows. In Section 2 we collect some basic facts on subnormalisers of semisimple elements lying in abelian Sylow subgroups of finite groups of Lie type. In Section 3 we determine in Theorem 3.1 the maximal overgroups of normalisers of Sylow $d$-tori in exceptional groups of simply connected Lie type and use this to describe in Theorem 3.2 the subnormalisers of semisimple $p$-elements in these groups in the case that Sylow $p$-subgroups are abelian. In Section 4 we solve the same problems for the various series of classical groups of Lie type, see in particular Theorems 4.6, 4.10, 4.17, 4.18, 4.25 and 4.27. In Sections 5 and 6 we complement our results by describing the subnormalisers of $p$-elements in symmetric as well as in sporadic simple groups with abelian Sylow $p$-subgroups and thus complete the proof of Theorem 2.
2. Subnormalisers of semisimple elements in the abelian Sylow case
Let $\mathbf{G}$} be a simple algebraic group of simply connected type with a Frobenius endomorphism $F$ with respect to an ${\mathbb{F}}_{q}$-structure and let $\ell$ be a prime. When Sylow $\ell$-subgroups of $G:={\mathbf{G}}^{F}$ are abelian the determination of subnormalisers relies on the knowledge of possible overgroups of normalisers of Sylow $d$-tori of $\mathbf{G}$}. See [18, §24] or [9, §3.5] for background on Sylow tori.
Proposition 2.1.
Let $\ell$ be a prime not dividing $q$ such that Sylow $\ell$-subgroups of $G={\mathbf{G}}^{F}$ are abelian. Let $d:=e_{\ell}(q)$ be the order of $q$ modulo $\ell$ and ${\mathbf{S}}_{d}$ be a Sylow $d$-torus of $\mathbf{G}$}. Then we have
  1. (a)
    $\ell>2$, $\ell$ is good for $\mathbf{G}$} and not a torsion prime;
  2. (b)
    $\Phi_{d}$ is the unique cyclotomic polynomial dividing the generic order of $({\mathbf{G}},F)$ with $\ell|\Phi_{d}(q)$;
  3. (c)
    ${\mathbf{S}}_{d}^{F}$ contains a Sylow $\ell$-subgroup $P$ of $G$; and
  4. (d)
    ${\operatorname{Sub}}_{G}(x)=\langle{\mathbf{C}}_{G}(x),{\mathbf{N}}_{G}({\mathbf{S}}_{d})\rangle$ for any $x\in P$.
Proof.
Since ${\operatorname{SL}}_{2}(q)$ and ${\operatorname{PGL}}_{2}(q)$ have non-abelian Sylow 2-subgroups, we must have $\ell>2$. By [15, Prop. 2.2], we have (b). Hence, any Sylow $d$-torus has order divisible by the full $\ell$-part of the order of $G$, giving (c). By inspection of the order formulae [18, Tab. 24.1], (b) also implies that $\ell$ does not divide the order of the Weyl group of $\mathbf{G}$} and thus is good for $\mathbf{G}$} and not a torsion prime (see [18, Tab. 14.1]), so we have (a). Finally, as $P$ is characteristic in ${\mathbf{S}}_{d}^{F}$, we have ${\mathbf{N}}_{G}({\mathbf{S}}_{d})\leq{\mathbf{N}}_{G}({\mathbf{S}}_{d}^{F})\leq{\mathbf{N}}_{G}(P)$; since $P$ is abelian, it is contained in a unique Sylow $d$-torus of $\mathbf{G}$} by [5, Prop. 2.2] and thus in fact ${\mathbf{N}}_{G}({\mathbf{S}}_{d})={\mathbf{N}}_{G}(P)$. Now (d) follows as ${\operatorname{Sub}}_{G}(x)=\langle{\mathbf{C}}_{G}(x),{\mathbf{N}}_{G}(P)\rangle$ by [16, Prop. 2.11]. ∎
We now dispose of the rank 1 case, that is, when Sylow $\ell$-subgroups of $G$ are cyclic:
Proposition 2.2.
[…]
Proof.
Let $1\neq x\in{\mathbf{S}}_{d}^{F}$ be an $\ell$-element. Since $\ell$ is good for $\mathbf{G}$} and not a torsion prime (see above), ${\mathbf{C}}_{\mathbf{G}}(x)$ is a $d$-split Levi subgroup of $\mathbf{G}$} by [5, Prop. 2.2]. Since $\Phi_{d}$ divides the generic order of $({\mathbf{G}},F)$ exactly once, the only $d$-split Levi properly containing ${\mathbf{C}}_{\mathbf{G}}({\mathbf{S}}_{d})$ is $\mathbf{G}$} itself. As we assumed $\mathbf{G}$} to be simple and $\ell$ is not a torsion prime, we have $x\notin{\mathbf{Z}}(G)$, so this forces ${\mathbf{C}}_{\mathbf{G}}(x)={\mathbf{C}}_{\mathbf{G}}({\mathbf{S}}_{d})$, whence we conclude by Proposition 2.1. ∎
We may and will hence assume in the sequel that $\Phi_{d}$ divides the order polynomial of our group at least twice and thus that Sylow $\ell$-subgroups have rank at least 2 (by [18, Thm 25.14]). In the following two sections we discuss the exceptional and the classical groups.
Remark 2.3.
Subnormalisers of $\ell$-elements of a given finite permutation group or matrix group over a finite field can be determined effectively using the criterion in [16, Cor. 2.10], which we have implemented in the GAP system [23]. This will be used throughout to treat small cases.
3. Overgroups of Sylow tori normalisers and subnormalisers in groups of exceptional type
Throughout this section, $\mathbf{G}$} is a simple linear algebraic group of simply connected type and $F:{\mathbf{G}}\to{\mathbf{G}}$ a Frobenius morphism with respect to an ${\mathbb{F}}_{q}$-structure. (Subnormalisers in the Suzuki and Ree groups were already handled in [16, Thm 5.11].) We consider primes $\ell$ not dividing $q$. Since subnormalisers of $\ell$-elements contain the normaliser of a Sylow $\ell$-subgroup, which in our case, agrees with the normaliser of a Sylow $d$-torus, we first classify overgroups of the latter type of subgroups.
Theorem 3.1.
Assume that $G:={\mathbf{G}}^{F}$ is of exceptional type, let $d\geq 1$ and let ${\mathbf{S}}_{d}\leq{\mathbf{G}}$ be a Sylow $d$-torus. Let $\ell$ be a prime with $d=e_{\ell}(q)$ such that Sylow $\ell$-subgroups of $G$ are abelian of rank at least $2$. Then ${\mathbf{N}}_{G}({\mathbf{S}}_{d})$ is maximal in the cases listed in Table 1, while the proper overgroups of ${\mathbf{N}}_{G}({\mathbf{S}}_{d})$ in the other cases are as given in Tables 2 and 3.
The last column in Tables 2 and 3, headed ${\mathbf{C}}_{G}(x)$, will be explained in Theorem 3.2.
Table 1. Maximal Sylow $d$-torus normalisers
[…]
Here, Sylow tori are indicated by their order polynomial, $W(E_{i})$ denotes a Weyl group of type $E_{i}$, and $G_{i}$ denotes a primitive complex reflection group according to Shephard–Todd.
Table 2. Generic overgroups of Sylow $d$-torus normalisers
[…]
Table 3. Non-generic overgroups of Sylow $d$-torus normalisers
$\begin{array}[]{ccccc}{\mathbf{G}}^{F}&d&{\mathbf{N}}_{G}({\mathbf{S}}_{d})&\text{overgroups}&{\mathbf{C}}_{G}(x)\\ \hline\cr G_{2}(4)&2&5^{2}.W(G_{2})&J_{2}&5.A_{1}(4)\\ F_{4}(2)&4&5^{2}.G_{8}&{}^{2}\!F_{4}(2)&\text{--}\\ F_{4}(3)&6&7^{2}.G_{5}&{}^{3}\!D_{4}(2).3&\text{--}\\ E_{6}(2)&4&5^{2}.G_{8}&{}^{2}\!F_{4}(2),F_{4}(2)&\text{--}\\ {}^{2}\!E_{6}(2)&3&7^{2}.3.G_{5}&F_{4}(2)\times 3$$\text{--}\\ \end{array}$
Proof.
We use knowledge on maximal subgroups of the groups $G$ in question. Note that $\ell>3$ since Sylow 2- and 3-subgroups of all groups considered are non-abelian. Thus by our assumptions that $d=e_{\ell}(q)$, $\ell$ is a Zsigmondy prime divisor of $\Phi_{d}(q)$, and ${\mathbf{N}}_{G}({\mathbf{S}}_{d})$ contains an abelian Sylow $\ell$-subgroup of $G$, by [18, Thm 25.14], of rank at least 2. The normalisers of Sylow $d$-tori are given in [3, Tab. 1 and 3]. For $G=G_{2}(q)$ it follows from the lists in [2, Tab. 8.30, 8.41 and 8.42] (for $p=2$, $p\geq 5$ and $p=3$, respectively) that the only maximal subgroups that can possibly contain ${\mathbf{N}}_{G}({\mathbf{S}}_{d})$ are ${\operatorname{SL}}_{3}(\pm q).2$, and in addition $J_{2}$ for $q=4$ and $d=2$. By [2, Tab. 8.3 and 8.5] there are no other proper subgroups of ${\operatorname{SL}}_{3}(\pm q).2$ containing ${\mathbf{N}}_{G}({\mathbf{S}}_{d})$. Here note that $q\geq 8$ for $d=1$ and $q\neq 2,3,5,7$ for $d=2$ as otherwise there is no prime $\ell$ as required.
For $G={}^{3}\!D_{4}(q)$, the result of Kleidman cited in [2, Tab. 8.51] shows that the only maximal overgroups of ${\mathbf{N}}_{G}({\mathbf{S}}_{d})$ for $d=1,2$ are as listed while for $d=3,6$, ${\mathbf{N}}_{G}({\mathbf{S}}_{d})$ is maximal. Again we conclude with [2, Tab. 8.3–8.6] that the only possible overgroups for $d=1,2$ are the stated ones.
For $G=F_{4}(q)$, the results of Craven [8, Tab. 1, 7 and 8] yield the given maximal subgroups as possible overgroups. Again by [2, Tab. 8.50], respectively our previous result for $^{3}\!D_{4}(q)$, there are no further non-maximal overgroups.
For $G=E_{6}(q)$ the tables [8, Tab. 2 and 9] allow us to conclude, and for $G={}^{2}\!E_{6}(q)$, the tables in [8, Tab. 3 and 10]. For types $E_{7}$ and $E_{8}$ we use [13, Thm 0] in conjunction with [7, Thm 1.1 and 1.2] to conclude that ${\mathbf{N}}_{G}({\mathbf{S}}_{d})$ is maximal in all cases listed in Table 1, and to derive the possible overgroups in the other cases, using also [18, Thm 29.1]. Observe that by assumption these need to contain an elementary abelian $\ell$-group, with $\ell>5$, of rank at least 2. ∎
Let us point out that the generic overgroups in Table 2 are all obtained as $F$-fixed points of a reductive subgroup $\mathbf{H}$} of $\mathbf{G}$} as follows: the connected component $\mathbf{H}$$^{\circ}$ is generated by the minimal $d$-split Levi subgroups properly containing ${\mathbf{N}}_{G}({\mathbf{S}}_{d})$ corresponding to one conjugacy class $C$ of reflections in the relative Weyl group $W_{d}$ of ${\mathbf{S}}_{d}$, and $({\mathbf{H}}/{\mathbf{H}}^{\circ})^{F}\cong W_{d}/\langle C\rangle$. This is the exact analogue of the corresponding result for subnormalisers of semisimple elements in simple algebraic groups in [16, Thm 6.8]. Nevertheless, the additional overgroups in Table 3 for small values of the parameters indicate that there may not be any conceptual proof of the preceding classification, avoiding the precise knowledge of all maximal subgroups.
Theorem 3.2.
Let $\mathbf{G}$} be simple with Frobenius map $F$ with respect to an ${\mathbb{F}}_{q}$-structure such that $G={\mathbf{G}}^{F}$ is of exceptional type. Let $\ell$ be a prime dividing $|G|$ such that Sylow $\ell$-subgroups of $G$ are abelian. Let $1\neq x\in G$ be an $\ell$-element and ${\mathbf{S}}_{d}\leq{\mathbf{G}}$ be a Sylow $d$-torus containing $x$, where $d=e_{\ell}(q)$. Then ${\operatorname{Sub}}_{G}(x)=G$ unless one of:
  1. (1)
    ${\mathbf{C}}_{G}(x)={\mathbf{C}}_{G}({\mathbf{S}}_{d})$, $x$ is picky and ${\operatorname{Sub}}_{G}(x)={\mathbf{N}}_{G}({\mathbf{S}}_{d})$; or
  2. (2)
    ${\mathbf{C}}_{G}(x)>{\mathbf{C}}_{G}({\mathbf{S}}_{d})$ is the $d$-split Levi subgroup in the last column of Tables 2 or 3, and ${\operatorname{Sub}}_{G}(x)$ is as in the next to last column.
Proof.
Since Sylow $\ell$-subgroups of $G$ are supposed to be abelian, we have ${\operatorname{Sub}}_{G}(x)=\langle{\mathbf{C}}_{G}(x),{\mathbf{N}}_{G}(P)\rangle$ by Proposition 2.1. By [16, Thm 5.3] the element $x$ is picky if and only if we are in Case (1), and then ${\operatorname{Sub}}_{G}(x)$ is as stated. Now assume ${\mathbf{C}}_{G}(x)$ properly contains ${\mathbf{C}}_{G}({\mathbf{S}}_{d})$. Since $\ell$ is a good prime for $\mathbf{G}$}, then ${\mathbf{C}}_{\mathbf{G}}(x)$ is a $d$-split Levi subgroup of $\mathbf{G}$}, see [5, Prop. 2.2], properly containing ${\mathbf{C}}_{\mathbf{G}}({\mathbf{S}}_{d})$. The possibilities can be obtained in Chevie [21] and are also listed in [9, Tab. 3.3] (unfortunately with some omissions for $E_{8}$ when $d=8,12$). By [16, Prop. 5.2], ${\operatorname{Sub}}_{G}(x)$ is generated by the $F$-fixed points of the normalisers of the Sylow $d$-tori of $\mathbf{G}$} containing $x$. Thus, either ${\operatorname{Sub}}_{G}(x)=G$, or it is one of the groups in Theorem 3.1. In the four cases with an entry “–” in the last column in Table 2, the stated overgroup does not contain any $d$-split Levi subgroup of $\mathbf{G}$} properly containing ${\mathbf{N}}_{G}({\mathbf{S}}_{d})$, so here ${\operatorname{Sub}}_{G}(x)=G$. In all other cases, the listed $d$-split Levi is the unique one embedding in the listed overgroup. In Table 3 it is easy to see from the known character tables and information on maximal subgroups that only in the stated case the listed overgroup of ${\mathbf{N}}_{G}({\mathbf{S}}_{d})$ does contain the centraliser of an $\ell$-element $x$ with ${\mathbf{C}}_{G}(x)>{\mathbf{C}}_{G}({\mathbf{S}}_{d})$. Note that the prime $\ell$ is uniquely determined in each of the cases in Table 3, namely $\ell=5,5,7,5,7$ in the respective cases. In $G_{2}(4)$, a GAP-computation shows that both the generic case ${\operatorname{SU}}_{3}(4).2$ and the exotic case $J_{2}$ do occur as subnormalisers for suitable 5-elements. ∎
We can check Conjecture 1 for the exceptional subnormaliser:
Proposition 3.3.
Conjecture 1 holds for $G_{2}(4)$ at $\ell=5$ with ${\operatorname{Sub}}_{G}(x)=J_{2}$.
Proof.
From the known character tables one sees that both $G$ and ${\operatorname{Sub}}_{G}(x)=J_{2}$ possess 14 irreducible character of degree prime to 5 not vanishing on $x$, and four further ones of degree divisible by 5 exactly once, and there is a bijection such that the values at $x$ of corresponding characters agree up to sign, so in particular generate the same field extension. ∎
4. Overgroups of Sylow tori normalisers and subnormalisers in groups of classical type
Here, throughout we let $\mathbf{G}$} be simple of simply connected classical type and $F:{\mathbf{G}}\to{\mathbf{G}}$ a Frobenius map with respect to an ${\mathbb{F}}_{q}$-structure, not inducing triality. Let $G:={\mathbf{G}}^{F}$. Let $\ell>2$ be a prime not dividing $q$ and $H$ an elementary abelian $\ell$-subgroup of $G$ of maximal possible rank. As $\ell$ is not a torsion prime for $\mathbf{G}$}, $H$ embeds into a maximal torus $\mathbf{T}$} of $\mathbf{G}$} by [18, Cor. 14.17]. Thus, $\mathbf{T}$} lies in the $F$-stable subgroup ${\mathbf{C}}_{\mathbf{G}}(H)$, which possesses $F$-stable maximal tori, all of which contain $H$. Hence, without loss we may assume $\mathbf{T}$} to be $F$-stable. Furthermore, $\mathbf{T}$} then contains a Sylow $d$-torus ${\mathbf{S}}_{d}$ of $({\mathbf{G}},F)$ for $d=e_{\ell}(q)$, by [14, Prop. 5.7], and $H\leq{\mathbf{S}}_{d}$. Thus $H$ is normalised by $N:={\mathbf{N}}_{G}({\mathbf{S}}_{d})$. We are thus led to studying overgroups of ${\mathbf{N}}_{G}({\mathbf{S}}_{d})$. This we will do in terms of Aschbacher’s classification, by which any maximal subgroup either lies in a collection ${\mathcal{C}}_{i}$ of natural, geometric subgroups, or else is almost simple modulo its centre and lies in a class denoted $\mathcal{S}$} (see [12], [2], or [18, §27] for an introduction and further references).
In all types we denote by $P_{r}$ a maximal parabolic subgroup of $G$ stabilising an $r$-dimensional totally singular subspace of the natural ${\mathbb{F}}_{q}$- respectively ${\mathbb{F}}_{q^{2}}$-module.
4.1. The special linear groups
We start our investigation with the groups ${\operatorname{SL}}_{n}(q)$. We fix the following setup and notation throughout this subsection. Let $G={\operatorname{SL}}_{n}(q)$ with $n\geq 2$, $(n,q)\neq(2,2),(2,3)$. Let $d\geq 1$ be an integer and write $n=ad+r$ with $0\leq r<d$. Let ${\mathbf{S}}_{d}<{\mathbf{G}}={\operatorname{SL}}_{n}$ be a Sylow $d$-torus. To avoid certain degenerate situations and since this will be satisfied for our intended application, we also assume that there is some Zsigmondy primitive prime divisor $\ell>2$ of $q^{d}-1$. So, in particular, $\ell$ divides $|G|$ if $d\leq n$. We first consider maximal subgroups of $G$ of geometric type (in Aschbacher’s approach).
Proposition 4.1.
Let $G={\operatorname{SL}}_{n}(q)$, and $d,\ell$ as above. If $M<G$ is a maximal subgroup containing the normaliser $N:={\mathbf{N}}_{G}({\mathbf{S}}_{d})$ of a Sylow $d$-torus ${\mathbf{S}}_{d}\leq{\mathbf{G}}$, then one of:
  1. (1)
    $M=P_{ad}$ or $P_{r}$ is maximal parabolic, if $r>0$;
  2. (2)
    $M=({\operatorname{GL}}_{d}(q)\wr{\mathfrak{S}}_{a})\cap G$ if $n=ad$ and $a>1$;
  3. (3)
    $M={\operatorname{GL}}_{n/t}(q^{t}).t\cap G$ if $n=d$ and $t|n$ is a prime;
  4. (4)
    $M={\operatorname{Sp}}_{n}(2)$ if $n>2$ is even and $q=d=2$;
  5. (5)
    $M={\operatorname{SU}}_{n}(2)$ if $n\geq 3$, $q=4$ and $d=1$; or
  6. (6)
    $M$ is in class $\mathcal{S}$}.
Note that in cases (4) and (5) we necessarily have $\ell=3$.
Proof.
By [9, Exmp. 3.5.14] the centraliser and normaliser of ${\mathbf{S}}_{d}$ have the structure
${\mathbf{C}}_{G}({\mathbf{S}}_{d})=\big({\operatorname{GL}}_{1}(q^{d})^{a}\times{\operatorname{GL}}_{r}(q)\big)\cap G\quad\text{and}\quad N=\big({\operatorname{GL}}_{1}(q^{d})\wr G(d,1,a)\times{\operatorname{GL}}_{r}(q)\big)\cap G$
(all viewed as subgroups of ${\operatorname{GL}}_{n}(q)$). Observe that for $d=1$, NN contains a subgroup $N_{0}:=C_{\ell}^{n-1}\rtimes{\mathfrak{S}}_{n}$, where the base group is the deleted permutation module for the symmetric group ${\mathfrak{S}}_{n}$, while for $d>1$ it contains a subgroup $N_{0}:=C_{\ell}^{a}\rtimes G(d,1,a)$ where here the complement acts irreducibly on the base group. In either case, the $N$-composition factors of the natural module have dimensions $ad$ and $r$, and $N$ acts primitively on the factor of dimension $r$. For the proof we now go through the various Aschbacher classes of maximal subgroups of $G$ as described, for example, in [18, Prop. 28.1].
Assume first $M=P_{m}$, with $1\leq m<n$, is a maximal parabolic subgroup with $N\leq M$, so $M$ acts maximally reducibly. Then by what we just observed, $m\in\{ad,r\}$, and $P_{m}$ is proper if $r>0$, so we arrive at case (1).
Next assume $M=({\operatorname{GL}}_{m}(q)\wr{\mathfrak{S}}_{t})\cap G$, with $n=mt$, $t\geq 2$, is imprimitive. If $d=1$ then the automiser in $M$ of an elementary abelian $\ell$-subgroup $E$ of ${\operatorname{GL}}_{m}(q)^{t}\cap G$ of rank n1n-1 is ${\mathfrak{S}}_{m}\wr{\mathfrak{S}}_{t}$, a proper subgroup of the automiser ${\mathfrak{S}}_{n}$ in $G$, unless $m=1$ as in case (2). If $d>1$ write $m=a_{1}d+r_{1}$ with $0\leq r_{1}<d$. Then $M$ has $\ell$-rank $ta_{1}$, so we need $ta_{1}\geq a$ whence $ta_{1}=a$. Then the automiser in $M$ of $E$ is $G(d,1,a_{1})\wr{\mathfrak{S}}_{t}$, again a proper subgroup of $G(d,1,a)$ unless $a_{1}=1$, so $t=a$ and hence $r=ar_{1}$. For its centraliser in $M$ to contain a subgroup ${\operatorname{GL}}_{r}(q)$ we need $r=0$ and so again obtain (2).
If $M={\operatorname{GL}}_{m}(q^{t}).t\cap G$, with $n=mt$, $t$ prime, is an extension field subgroup, then $M$ contains an elementary abelian $\ell$-group of rank $a$, respectively of rank n1n-1 if $d=1$, only if $t|\gcd(r,d)$. Its automiser in $M$ is then $G(d/t,1,a).t$, which equals $G(d,1,a)$ only if $a=1$. Comparing the centralisers of a Sylow $d$-torus in $M$ and $G$ we then see that $r=0$. So we must have $n=d$, giving (3).
Next assume $M={\operatorname{GL}}_{n_{1}}(q)\otimes{\operatorname{GL}}_{n_{2}}(q)\cap G$ with $n_{1}n_{2}=n$, $1<n_{1}<n_{2}$, preserves a tensor decomposition. Writing $n_{i}=a_{i}d+r_{i}$ with $0\leq r_{i}<d$, the $\ell$-rank of $M$ is $a_{1}+a_{2}$, while $G$ has $\ell$-rank at least $a_{1}a_{2}d+a_{1}r_{2}+a_{2}r_{1}-1$, and this is bigger unless $r_{1}=r_{2}=0$ and either $a_{1}=a_{2}=1$, $d=2$, so $n_{1}=n_{2}=2$ which is excluded, or $a_{1}=2,a_{2}=3$, $d=1$. In the latter case the automiser in $G$ is ${\mathfrak{S}}_{6}$, while it is only ${\mathfrak{S}}_{2}\times{\mathfrak{S}}_{3}$ in $M$. The same reasoning applies to $M={\operatorname{GL}}_{m}(q)\wr{\mathfrak{S}}_{t}$ with $n=m^{t}$, $t\geq 2$, $m\geq 3$.
Now let $M=t^{2m+1}.{\operatorname{Sp}}_{2m}(t){\mathbf{Z}}(G)$ be the normaliser of an extra-special $t$-subgroup with $n=t^{m}$, where furthermore $q$ is minimal among powers of $p$ with $q\equiv 1\pmod{t(2,t)}$. The maximal order of an abelian subgroup in $M$ is $t^{m(m+1)/2+m+1}\gcd(n,q-1)$ by [25, Thm 3.1], while $N$ contains a maximal torus of order at least $(q-1)^{n-1}$. The ensuing inequality does not hold for any $n\geq 5$.
For $M={\operatorname{Sp}}_{n}(q)$ with $n$ even to have large enough $\ell$-rank, the order formulas show that we need $d$ to be even. In this case, $M$ contains a subgroup isomorphic to $N_{0}$, in the normaliser of a Sylow $d$-torus of $M$. But the semisimple part of its centraliser in $M$ is a symplectic group ${\operatorname{Sp}}_{r}(q)$, properly smaller than ${\operatorname{GL}}_{r}(q)$ unless $r=0$ (note that here $r$ is necessarily even as both $n,d$ are). In the latter case, ${\mathbf{C}}_{G}({\mathbf{S}}_{d})$ contains a subgroup of index q1q-1 of a homocyclic group $(q^{d}-1)^{a}$, with $n=ad$, and such an abelian subgroup can only be contained in $M$ if $q^{d/2}-1=1$, so $q=d=2$ (and thus $\ell=3$) as in (4).
If $M={\operatorname{SO}}_{n}^{(\pm)}(q)$ with $q$ odd, arguing as in the previous case we reach the contradiction $q=2$. For $M={\operatorname{SU}}_{n}(q_{0})$ with $q=q_{0}^{2}$ the same type of consideration leads to possibility (5).
Finally, assume $M={\operatorname{GL}}_{n}(q_{0})\cap G$, with $q=q_{0}^{f}$, $f\geq 2$, is a subfield group. Note that $N$ contains a maximal torus of $G$, of order at least $(q^{d}-1)^{a}(q-1)^{n-da-1}$, while the size of maximal tori in $M$ is bounded above by $(q_{0}+1)^{n-1}$, with $q_{0}^{2}\leq q$. The resulting inequality forces $q=4$, $q_{0}=2$ and thus $\ell=3$, $d=1$, but in this case the $3$-rank of $M$ is at most $\lfloor(n+1)/2\rfloor$ while $N_{0}$ has $3$-rank n1n-1, forcing $n\leq 3$. But for $n=2$, $M$ acts imprimitively and already occurs under (2), while for $n=3$ it does not contain a subgroup $N_{0}$. ∎
In order to deal with the maximal subgroups in class $\mathcal{S}$}, we first derive some easy estimates.
Lemma 4.2.
Let $n\geq 2$ and $q$ be a prime power. Then $|{\operatorname{SL}}_{n}(q)|>(q-1)q^{n^{2}-2}$.
Proof.
By Euler’s pentagonal number theorem, for all real $x$ with $|x|<1$ we have
$\prod_{i=1}^{\infty}(1-x^{i})=1+\sum_{k=1}^{\infty}(-1)^{k}\big(x^{k(3k+1)/2}+x^{k(3k-1)/2}\big).$
In particular, if $0\leq x\leq 1/2$ then
$\prod_{i=1}^{\infty}(1-x^{i})=1-x-x^{2}+x^{5}+x^{7}\ldots\geq 1-x-x^{2}+x^{5}=(1-x)(1-x^{2}-x^{3}-x^{4}).$
Thus, for $x:=q^{-1}$ and $n\geq 2$ we obtain
$\prod_{i=2}^{n}\frac{q^{i}-1}{q^{i}}=\prod_{i=2}^{n}(1-q^{-i})\geq 1-q^{-2}-q^{-3}-q^{-4}>\frac{q-1}{q},$
showing that
$\prod_{i=2}^{n}(q^{i}-1)>(q-1)q^{\binom{n+1}{2}-2}.$
Our claim is immediate from this with the order formula for ${\operatorname{SL}}_{n}(q)$. ∎
Lemma 4.3.
The subgroup $N$ of $G={\operatorname{SL}}_{n}(q)$ has order $|N|\geq q^{n-1}$.
Proof.
By our assumptions we have $d>1$ if $q\leq 3$, and thus $q^{d}-1\geq 3q^{d}/4$. Then since $|G(d,1,a)|=d^{a}|{\mathfrak{S}}_{a}|=d^{a}a!\geq(4/3)^{a}$ we conclude that $(q^{d}-1)^{a}|G(d,1,a)|\geq q^{ad}$ unless $ad=1$. But $ad>1$ in our situation. Since $|N|=(q^{d}-1)^{a}|G(d,1,a)|\cdot|{\operatorname{SL}}_{r}(q)|$, our claim follows from Lemma 4.2 when $r\geq 2$. Direct computation shows that it also holds when $r\in\{0,1\}$. ∎
Proposition 4.4.
In the situation of Proposition 4.1, assume $M$ is a maximal subgroup of $G$ in class $\mathcal{S}$} containing $N$. Then $M<G$ is one of:
  1. (1)
    $2.{\mathfrak{A}}_{5}<{\operatorname{SL}}_{2}(9)$, with $d=2$, $\ell=5$;
  2. (2)
    $2.{\mathfrak{A}}_{5}<{\operatorname{SL}}_{2}(11)$, with $d=1$, $\ell=5$;
  3. (3)
    $3.{\mathfrak{A}}_{6}<{\operatorname{SL}}_{3}(4)$, with $d=1$, $\ell=3$ or $d=2$, $\ell=5$; or
  4. (4)
    ${\mathfrak{A}}_{7}<{\operatorname{SL}}_{4}(2)$, with $d=3$, $\ell=7$.
Proof.
We consider the various possibilities for the non-abelian simple composition factor $S:=F^{*}(M)/{\mathbf{Z}}(F^{*}(M))$ of $M$ according to the classification. First assume $S$ is not of Lie type in the same characteristic as $G$.
(1) For $n\leq 12$ the possible $M\in{\mathcal{S}}$ are given in the tables of [2, §8]. By Lemma 4.3, if $N\leq M$ we must have
$|{\operatorname{Aut}}(S)|\geq|N|/|{\mathbf{Z}}(G)|\geq q^{n-1}/|{\mathbf{Z}}(G)|\geq q^{n-1}/n,$
which for any $S$ gives a small upper bound on $q$. Also, $\bar{N}:=N{\mathbf{Z}}(G)/{\mathbf{Z}}(G)$ contains elements of order at least q1q-1, so ${\operatorname{Aut}}(S)$ must contain elements of at least that order, yielding further restrictions on $q$. The occurring subgroups $M\in{\mathcal{S}}$ can now be investigated using the Atlas [6], leading to the four items listed in the conclusion. Hence from now on we can assume $n\geq 13$. Then $N$ contains a maximal torus of $G$ of size at least $(q^{2}-1)^{6}\geq 3^{6}$ when $q\leq 3$, respectively $(q-1)^{11}\geq 3^{11}$ when $q\geq 4$,
(2) Now assume $S={\mathfrak{A}}_{m}$ is an alternating group. If $S$ has a faithful projective representation of degree less than m2m-2 then $m\leq 8$ by [12, Prop. 5.3.7] so since we have $n\geq 13$, we may assume $m\leq n+2$. As argued in the proof of Lemma 4.3 then $q^{d}-1\geq 3q^{d}/4$. If $r\geq 2$ then $\bar{N}$ contains an abelian subgroup of size $(q^{d}-1)^{a}$, and $(q^{d}-1)^{a}\geq(4^{d}-1)^{a}\geq(4-1)^{ad}=3^{ad}$ if $q\geq 4$. But the maximal size of an abelian subgroup of ${\mathfrak{S}}_{m}$ is bounded above by $3^{m/3}$ by [4, Thm 1], and since $ad>n/2$ we conclude we must have
$3^{n/2}<3^{ad}<3^{(n+2)/3}$
whence $n<4$, a contradiction. If $q=3$ the existence of $\ell$ forces $d\geq 3$ and then $3^{d}-1\geq 3^{5d/6}$. So in this case we deduce $(3^{d}-1)^{a}\geq 3^{5ad/6}$ must be less than $3^{5n/12}\leq 3^{(n+2)/3}$, whence $n<8$. If $r\leq 1$ and still $q\geq 3$ then $\bar{N}$ will still contain an abelian subgroup of size at least $(q^{d}-1)^{a}/n(q-1)$, but we also have $ad\geq n-1$ and again the required inequality is not satisfied for $n>11$.
It remains to consider the possibility that $q=2$. First assume $r\leq 2$. As $d\geq 2$ we have $2^{d}-1\geq 2^{4d/5}$ and so conclude we must have $2^{4(n-2)/5}\leq 3^{(n+2)/3}$, so 4(n2)/5<8(n+2)/154(n-2)/5<8(n+2)/15 which implies $n<10$, a case already considered. Finally, if $r>2$ then we use that $N$ contains an elementary abelian $\ell$-group of rank $a$ centralised by a subgroup ${\operatorname{SL}}_{r}(2)$. Now the centraliser in ${\mathfrak{S}}_{m}$ of such an elementary abelian $\ell$-subgroup is ${\mathfrak{S}}_{m-a\ell}$ times an $\ell$-group, so ${\operatorname{SL}}_{r}(2)$ must embed into ${\mathfrak{S}}_{m-a\ell}$ and hence possess a faithful representation in characteristic 0 of degree at most $m-a\ell-1$. This implies $2^{r-1}\leq m-a\ell-1$ by [24]. On the other hand, as $\ell$ is a Zsigmondy prime divisor of $q^{d}-1$ we have $\ell\geq d+1$. So
$2^{r-1}\leq m-a\ell-1\leq n+2-a(d+1)-1=1-a+r\leq r$
which is never satisfied for $r>2$.
(3) Next assume $S$ is of Lie type in characteristic not dividing $q$. For $S={\operatorname{L}}_{m}(y)$ with $m\geq 3$ the smallest degree of a faithful projective representation in characteristic not dividing $y$ is $(y^{m}-1)/(y-1)-m\geq y^{m-1}$ by [24, Tab. 1], unless $(m,y)\in{\mathcal{E}}:=\{(3,2),(3,4),(4,2),(4,3)\}$. Furthermore,
$|M|\leq|{\operatorname{Aut}}(S)|\leq 2|{\operatorname{PGL}}_{m}(y)|\log_{2}y\leq 2\ y^{m^{2}-1}\log_{2}y.$
Thus, for $(m,y)\notin{\mathcal{E}}$, if $N\leq M$ then
$q^{y^{m-1}-1}\leq q^{n-1}\leq|N|\leq|M|\leq 2y^{m^{2}-1}\log_{2}y$
by Lemma 4.3. This is only satisfied for $(m,y)=(3,3)$. It thus remains to discuss $(m,y)\in{\mathcal{E}}\cup\{(3,3)\}$. All projective irreducible representations of ${\operatorname{L}}_{3}(2)$ have degree at most 8, the group ${\operatorname{L}}_{3}(3)$ does not have an irreducible representation of degree $n>12$ satisfying the inequality, for ${\operatorname{L}}_{3}(4)$ and ${\operatorname{L}}_{4}(2)\cong{\mathfrak{A}}_{8}$ the inequality only holds when $n\leq 12$, and for ${\operatorname{L}}_{3}(4)$ it is never satisfied. Thus no further cases arise.
For $S={\operatorname{L}}_{2}(y)$, $y\geq 7$ and $y\neq 9$ (as we already considered alternating groups), the normaliser of a (cyclic) Sylow $\ell$-subgroup has order at most y(y1)y(y-1), while the minimal projective degree is (y1)/(2,y1)(y-1)/(2,y-1), so we need
$y(y-1)\geq q^{n-1}\geq q^{(y-1)/2}-1.$
This is not satisfied for any $n\geq 13$ and prime powers $y,q$, so no further examples arise.
For $S={\operatorname{U}}_{m}(y)$ with $m\geq 3$ the smallest projective faithful degree is at least $(y-1)y^{m-2}$, unless $(m,y)\in\{(4,2),(4,3)\}$, by [24, Tab. I], and $|{\operatorname{Aut}}(S)|\leq 2y^{m^{2}-1}\log y$. Arguing as above, the only cases that satisfy the relevant inequality are
$(m,y)\in\{(3,3),(3,4),(3,5),(5,2),(6,2)\}.$
A maximal abelian subgroup of ${\operatorname{U}}_{3}(y)$ has size $(y+1)^{2}$ by [25, Thm 3.1]. But as pointed out above, for $n\geq 13$ there is a torus in $N$ of size at least $(q^{2}-1)^{6}$ when $q\leq 3$, respectively $3^{11}$ when $q\geq 4$, hence the case $m=3$ does not lead to new examples. For $S={\operatorname{U}}_{4}(2)$ or ${\operatorname{U}}_{4}(3)$, ${\operatorname{Aut}}(S)$ contains no abelian subgroups of the required order. For $S={\operatorname{U}}_{5}(2)$ the relevant inequality only holds when $q=3$ and $n\leq 16$, but by [11] the smallest degree $n>12$ of a 3-modular projective irreducible representation of $S$ is 44. For $S={\operatorname{U}}_{6}(2)$ the inequality only holds when $q=3$ and $n\leq 23$, while by [11] the only 3-modular projective irreducible representation of $S$ in degree $12<n\leq 23$ has degree 21, in which case $|{\mathbf{S}}_{d}^{F}|$ is too large.
For $S={\operatorname{S}}_{2m}(y)$ with $m\geq 2$ the smallest faithful projective degree is $(y^{m}-1)/2$ if $y$ is odd, and $(y^{m}-1)(y^{m}-y)/(2(y+1))$ if $y$ is even. (Note that we may assume $(m,y)\neq(2,2)$ as ${\operatorname{S}}_{4}(2)\cong{\mathfrak{S}}_{6}$.) Our inequality (with $n>12$) is satisfied only for $(2m,y)\in\{(4,3),(4,5),(4,7),(6,2),(6,3),(8,3)\}$. The condition that ${\operatorname{Aut}}(S)$ should contain an abelian subgroup of size $|{\mathbf{S}}_{d}^{F}|/{\mathbf{Z}}(G)$ rules out all but $S={\operatorname{S}}_{6}(3)$ with $q=2$. The smallest faithful irreducible 2-modular representations of $S$ have dimension 13 and 78. But those of degree 13 are only defined over ${\mathbb{F}}_{4}$, while the order of $N$ in ${\operatorname{SL}}_{78}(2)$ is much too large.
For $S={\operatorname{O}}_{2m+1}(y)$ with $m\geq 3$ and $y$ odd by [24, Tab. 1] the smallest projective degree is at least $(y^{m}-1)(y^{m}-y)/(y^{2}-1)$, unless $(m,y)=(3,3)$, and the necessary inequality is never satisfied. For $S={\operatorname{O}}_{7}(3)$ the smallest faithful projective degree is 27 by [11], but ${\operatorname{Aut}}(S)$ has no abelian subgroup of size at least $3^{12}$.
For $S={\operatorname{O}}_{2m}^{+}(y)$ with $m\geq 4$, again by [24] the minimal projective degree is at least $(y^{m}-1)(y^{m-1}-1)/(y^{2}-1)-7$, respectively 8 for ${\operatorname{O}}_{8}^{+}(2)$. Our inequality is never satisfied in the former case; for $S={\operatorname{O}}_{8}^{+}(2)$ we obtain $q=3$, but by [11] there is no projective irreducible 3-modular representation of $S$ of degree $n\geq 13$ for which the inequality would be satisfied.
For ${\operatorname{O}}_{2m}^{-}(q)$ with $m\geq 4$ the minimal degree is at least $(y^{m}+1)(y^{m-1}-y)/(y^{2}-1)-m+2$ and our inequality is never satisfied. For $S$ of exceptional Lie type, the lower bounds on faithful projective representations in [24] are always large enough to exclude these possibilities.
(4) Now assume $S$ is sporadic. Since $N$ and thus $M$ contain elements of order $(q^{d}-1)/(q-1)$ we see that $d\leq 6$, and $d\leq 5$ if $q\geq 3$. For $q=2$, $d=6$, there is no Zsigmondy prime, so we have $d\leq 5$ and hence $a\geq 2$ as $n\geq 13$. Thus $\ell^{2}$ divides $|S|$, forcing $\ell\leq 13$. By inspection, if $\ell^{a}$ divides $|{\operatorname{Aut}}(S)|$ then $n<(a+1)d\leq(a+1)(\ell-1)\leq 50$. Comparing to the list of degrees of faithful irreducible projective representations of $S$ below $50$ in [11] shows that no examples arise.
(5) Finally assume $S$ is of Lie type in the same characteristic as $G$. Again, when $n\leq 12$ the tables in [2, §8] show that no example exists, so assume $n\geq 13$. If $S={\operatorname{L}}_{m}(p^{f})$ then $n\geq m(m-1)/2$ and $p^{f}|q$, or $n\geq m^{k}$ and $p^{f}|q^{k}$ for some $k\geq 2$ by [12, Prop. 5.4.6, 5.4.11]. Now write $m=bd+s$ with $0\leq s<d$, so that $S$ has $\ell$-rank at most $b$. If $d=1$ then $G$ has $\ell$-rank at least $n-1\geq m(m-1)/2-1>m=b$, a contradiction, so $d\geq 2$. Then $n\geq m(m-1)/2=(bd+s)(bd+s-1)/2\geq d(b^{2}d/2+bs-b/2)$, so $G$ has $\ell$-rank at least $b^{2}d/2+bs-b/2>b$, again a contradiction unless $d=2,b=1,s=0$, but then $n\leq 12$.
If $S={\operatorname{S}}_{2m}(p^{f})$ then $n\geq 2m(2m-1)-2$ and $p^{f}|q$, or $n\geq(2m)^{k}$ and $p^{f}|q^{k}$ for some $k\geq 2$, or $m\leq 6$ and $n=2^{m}$, by [12, Tab. 5.4.A]. The first two cases are treated as before. To exclude $n=2^{m}$, with $4\leq m\leq 6$, observe that the $\ell$-rank of $S$ is at most $m/(d/2)$, and that of $G$ is at least $2^{m}/d-1>2m/d$. The same line of argument now applies to all types of groups $S$, using the bounds in [12, Tab. 5.4.A, 5.4.B]. For $S$ a triality group or of (possibly twisted) type $B_{2},G_{2}$ or $F_{4}$ we also refer to the description in [12, Rem. 5.4.7] for fields of definition. No further examples arise. ∎
Remark 4.5.
The previous results show, in particular, that in the situation of Proposition 4.1 the normaliser of a Sylow $d$-torus lies in a unique maximal subgroup of ${\operatorname{SL}}_{n}(q)$ whenever Sylow $\ell$-subgroups are non-cyclic and $r=0$, and $\ell\neq 3$ when $q\in\{2,4\}$. Note that in the latter exceptions, Sylow 3-subgroups of ${\operatorname{SL}}_{n}(q)$ are non-abelian. (Compare to the cases in [1, Tab. B].)
We can now determine the subnormalisers:
Theorem 4.6.
Let $G={\operatorname{SL}}_{n}(q)$ with $n\geq 2$, let $\ell\nmid q$ be a prime such that Sylow $\ell$-subgroups of $G$ are abelian and ${\mathbf{S}}_{d}\leq{\mathbf{G}}$ a Sylow $d$-torus, where $d=e_{\ell}(q)$. Write $n=ad+r$ with $0\leq r<d$. Then for $x\in{\mathbf{S}}_{d}^{F}$ an $\ell$-element we have
  1. (1)
    ${\operatorname{Sub}}_{G}(x)={\mathbf{N}}_{G}({\mathbf{S}}_{d})$ if ${\mathbf{C}}_{G}(x)={\mathbf{C}}_{G}({\mathbf{S}}_{d})$; or
  2. (2)
    ${\operatorname{Sub}}_{G}(x)=\big({\operatorname{GL}}_{ad}(q)\times{\operatorname{GL}}_{r}(q)\big)\cap G$ if $r>0$ and ${\mathbf{C}}_{{\operatorname{GL}}_{n}(q)}(x)=\prod_{i}{\operatorname{GL}}_{n_{i}}(q^{d})\times{\operatorname{GL}}_{r}(q)$ with $a=\sum n_{i}$ and at least one $n_{i}>1$; or
  3. (3)
    ${\operatorname{Sub}}_{G}(x)=\big({\operatorname{GL}}_{d}(q)\wr{\mathfrak{S}}_{a}\big)\cap G$ if $r=0$, $d>1$, $a>1$ and ${\mathbf{C}}_{{\operatorname{GL}}_{n}(q)}(x)={\operatorname{GL}}_{1}(q^{d})^{a-1}\times{\operatorname{GL}}_{d}(q)$; or
  4. (4)
    ${\operatorname{Sub}}_{G}(x)=G$ otherwise.
Proof.
By Proposition 2.1 the subnormaliser of $x$ is generated by $N:={\mathbf{N}}_{G}({\mathbf{S}}_{d})$ and ${\mathbf{C}}_{G}(x)$. Since ${\mathbf{G}}={\operatorname{SL}}_{n}$ is simply connected and Sylow $\ell$-subgroups of $G$ are abelian, the centraliser ${\mathbf{C}}_{\mathbf{G}}(x)$ is a $d$-split Levi subgroup of $({\mathbf{G}},F)$ by [5, Prop. 2.2]. First note that if $d>n/2$ then $\Phi_{d}$ divides the order polynomial of ${\operatorname{SL}}_{n}$ just once, and so ${\operatorname{Sub}}_{G}(x)={\mathbf{N}}_{G}({\mathbf{S}}_{d})$ and ${\mathbf{C}}_{G}(x)={\mathbf{C}}_{G}({\mathbf{S}}_{d})$ by Proposition 2.2 unless $x=1$, so we reach (1) or (4) of the conclusion. The same holds if ${\mathbf{C}}_{G}(x)={\mathbf{C}}_{G}({\mathbf{S}}_{d})\leq{\mathbf{N}}_{G}({\mathbf{S}}_{d})$ is a minimal $d$-split Levi.
If $d\leq n/2$ we go through the possible maximal overgroups of ${\mathbf{N}}_{G}({\mathbf{S}}_{d})$ classified in Propositions 4.1 and 4.4 to see which ones can possibly contain a non-minimal $d$-split Levi subgroup $L$ of $G$. By [9, Exmp. 3.5.14] the latter have the form
$\Big({\operatorname{GL}}_{n_{1}}(q^{d})\times\cdots\times{\operatorname{GL}}_{n_{t}}(q^{d})\times{\operatorname{GL}}_{s}(q)\Big)\cap G\qquad\text{with}\quad d\sum_{i}n_{i}+s=n$
(which of course implies $s\geq r$). As $L$ is non-minimal, we may assume that $s>r$ or $n_{1}>1$, say. First consider the overgroups in class $\mathcal{S}$} in Proposition 4.4. Only $M=3.{\mathfrak{A}}_{6}<{\operatorname{SL}}_{3}(4)$ with $d=1$, $\ell=3$ has non-cyclic Sylow $\ell$-subgroups. By explicit computation in GAP, this does not occur as a subnormaliser of a 3-element in ${\operatorname{SL}}_{3}(4)$.
Now consider the geometric subgroups of $G$ classified in Proposition 4.1, and first assume $r=0$. Then Cases (2)–(5) from Proposition 4.1 are relevant. In Case (3) we have $d=n$, so Sylow $\ell$-subgroups of $G$ are cyclic, a situation already discussed before. In Cases (4) and (5) we have $\ell=3$ and Sylow $\ell$-subgroups of $M$ are abelian and non-cyclic only for $M={\operatorname{Sp}}_{4}(2)<{\operatorname{SL}}_{4}(2)\cong{\mathfrak{A}}_{8}$, but direct computation shows that $M$ is not the subnormaliser of any 3-element of ${\operatorname{SL}}_{4}(2)$. Finally, in Case (2) the action on the natural module shows that for ${\mathbf{C}}_{G}(x)$ to be contained in $M$ we need the structure given in (3) of the conclusion. Here, since ${\operatorname{Sub}}_{G}(x)$ contains the ${\operatorname{GL}}_{d}(q)$-factor from ${\mathbf{C}}_{G}(x)$ as well as the symmetric group ${\mathfrak{S}}_{a}$ from the normaliser of a Sylow $d$-torus, we see that ${\operatorname{Sub}}_{G}(x)=M$, as claimed. Note that if $d=1$ we are in Case (1) and if $a=1$ we are in Case (4).
So, finally assume $r>0$. Then only the parabolic subgroups $M\in\{P_{ad},P_{r}\}$ occur in Proposition 4.1. Comparing dimensions of composition factors on the natural module we see that either can contain $L={\mathbf{C}}_{G}(x)$ as above only if $r=s$. Now note that the ${\operatorname{GL}}_{r}(q)$-factor times its centraliser in $G$ is conjugate to a Levi factor of $M$, and it contains the normaliser of a Sylow $d$-torus. Thus, if ${\operatorname{Sub}}_{G}(x)\leq M$ then it already lies in a Levi factor. Thus ${\operatorname{Sub}}_{G}(x)$ has the form $(H\times{\operatorname{GL}}_{r}(q))\cap G$ for some subgroup $H$ of ${\operatorname{GL}}_{ad}(q)$ containing a Sylow $d$-torus normaliser. So we can again appeal to Proposition 4.1 to see that if $H$ is proper, it must lie in a subgroup of type (2)–(6). Here note that the subgroups of type (2) act imprimitively on a sum of $a$ subspaces of dimension $d$, while $L$ has a primitive summand of dimension at least $2d$ (as $n_{1}>1$), so this case is out. In Case (3) the Sylow $\ell$-subgroups are cyclic, contrary to assumption, and Cases (4)–(6) do not occur by the same arguments as given in the previous paragraph. So here ${\operatorname{Sub}}_{G}(x)$ is a Levi factor of $M$, as in (1) of the conclusion. ∎
4.2. The special unitary groups
We next consider the special unitary groups. So throughout this subsection let $G={\operatorname{SU}}_{n}(q)$ with $n\geq 3$, $(n,q)\neq(3,2)$. Let $d\geq 1$ be an integer and
$e:=\begin{cases}2d&\text{if $d$ is odd},\\ d/2&\text{if $d\equiv 2\pmod{4}$},\\ d&\text{if $d\equiv 0\pmod{4}$.}\end{cases}$
Write $n=ae+r$ with $0\leq r<e$. We again assume there is a Zsigmondy primitive prime divisor $\ell>2$ of $q^{d}-1$; so we have $d=e_{\ell}(q)$ and $e=e_{\ell}(-q)$. Let ${\mathbf{S}}_{d}\leq{\mathbf{G}}$ be a Sylow $d$-torus with normaliser $N:={\mathbf{N}}_{G}({\mathbf{S}}_{d})$. By [9, Exmp. 3.5.14] we have
${\mathbf{C}}_{G}({\mathbf{S}}_{d})=\big({\operatorname{GL}}_{1}((-q)^{e})^{a}\times{\operatorname{GU}}_{r}(q)\big)\cap G\quad\text{and}\quad N=\big({\operatorname{GL}}_{1}((-q)^{e})\wr G(e,1,a)\times{\operatorname{GU}}_{r}(q)\big)\cap G.$
Proposition 4.7.
Let $G={\operatorname{SU}}_{n}(q)$ with $n\geq 3$ and $d,e,\ell$ as above. If $M<G$ is a maximal subgroup containing the normaliser $N:={\mathbf{N}}_{G}({\mathbf{S}}_{d})$ of a Sylow $d$-torus ${\mathbf{S}}_{d}\leq{\mathbf{G}}$, then one of:
  1. (1)
    $M=({\operatorname{GU}}_{ae}(q)\times{\operatorname{GU}}_{r}(q))\cap G$ if $r>0$;
  2. (2)
    $M=({\operatorname{GU}}_{e}(q)\wr{\mathfrak{S}}_{a})\cap G$ if $n=ae$ and $a>1$;
  3. (3)
    $M={\operatorname{GL}}_{n/2}(q^{2}).2\cap G$ if $n=e$ is even;
  4. (4)
    $M={\operatorname{GU}}_{n/t}(q^{t}).t\cap G$ if $n=e$ and $2<t|n$ is a prime;
  5. (5)
    $M={\operatorname{Sp}}_{4}(q).(q-1,2)$ if $n=d=4$ and $q\leq 3$; or
  6. (6)
    $M$ is in class $\mathcal{S}$}.
Proof.
The situation for ${\operatorname{SU}}_{n}(q)$ is Ennola dual to the one for ${\operatorname{SL}}_{n}(q)$. The arguments are now similar to the case of ${\operatorname{SL}}_{n}(q)$ in Proposition 4.1, where as far as divisibility questions are concerned, we need to replace $q$ by q-q and $d$ by $e$, and we now appeal to [12, Tab. 3.5B] for the description of the Aschbacher classes.
Assume first $M$ is reducible. Since the composition factors of $N$ on the natural module have dimensions $ae$ and $r$, we either have $M$ is as in (1), or $M$ is a parabolic subgroup $P_{r}$ with $r>0$ (note that $ae>n/2$ is larger than the dimension of a totally isotropic subspace). But the latter has three composition factors on the natural module for $G$.
The argument for the imprimitive groups $M=({\operatorname{GU}}_{m}(q)\wr{\mathfrak{S}}_{t})\cap G$ with $n=mt$ is identical to the one for ${\operatorname{SL}}_{n}(q)$ and we reach conclusion (2). The subgroups ${\operatorname{GL}}_{n/2}(q^{2}).2\cap G$ are normalisers of Levi subgroups of $G$ and it can be seen from the description of their Sylow $d$-normaliser in the proof of Proposition 4.1 that the automiser of a maximal rank $\ell$-subgroup is strictly smaller than $G(e,1,a)$ unless $a=1$, so $n=e$ as in (3).
By arguments as in the case of ${\operatorname{SL}}_{n}(q)$, the only further extension field subgroups that can occur are as in (4), while again there are no examples for stabilisers of tensor decompositions by order comparison. Let next $M=t^{1+2m}.{\operatorname{Sp}}_{2m}(t){\mathbf{Z}}(G)$ be the normaliser of an extra-special $t$-subgroup where $n=m^{t}$. As for the case of ${\operatorname{SL}}_{n}(q)$, a maximal abelian subgroup in $M$ has size at most $t^{m(m+1)/2+m+1}\gcd(n,q+1)$, while $N$ contains a maximal torus of order at least $(q-1)^{n-1}$, respectively $3^{(n-1)/2}$ if $q=2$. Comparing the orders we arrive only at the case $2^{1+8}.{\operatorname{Sp}}_{8}(2)<{\operatorname{SU}}_{16}(3)$. Here, the $\ell$-parts of $M$ and $G$ only agree for $\ell=17$, but then the centralisers of $\ell$-elements in $G$ are too large.
Similarly, by slight variations of the considerations for ${\operatorname{SL}}_{n}(q)$, none of the other types of geometric maximal subgroups apart from $M={\operatorname{Sp}}_{2n}(q)$ can contain $N$; for the latter we arrive at the condition $n=4=d$ for $M$ to contain an $\ell$-subgroup of sufficient rank, and furthermore $q\leq 3$ for it to contain the normaliser of a Sylow $d$-torus. ∎
Proposition 4.8.
In the situation of Proposition 4.7 assume $M$ is a maximal subgroup of $G$ in class $\mathcal{S}$} containing $N$. Then $M<G$ is one of:
  1. (1)
    ${\operatorname{L}}_{2}(7)<{\operatorname{SU}}_{3}(3)$, with $d=6$, $\ell=7$;
  2. (2)
    $3.{\mathfrak{A}}_{6}.2_{3}<{\operatorname{SU}}_{3}(5)$, with $d=2$, $\ell=3$;
  3. (3)
    $3.{\mathfrak{A}}_{7}<{\operatorname{SU}}_{3}(5)$, with $d=2$, $\ell=3$, or $d=6$, $\ell=7$;
  4. (4)
    $4\circ 2.{\mathfrak{A}}_{7}<{\operatorname{SU}}_{4}(3)$, with $d=4$, $\ell=5$, or $d=6$, $\ell=7$;
  5. (5)
    $4_{2}.{\operatorname{L}}_{3}(4)<{\operatorname{SU}}_{4}(3)$, with $d=6$, $\ell=7$;
  6. (6)
    ${\operatorname{L}}_{2}(11)<{\operatorname{SU}}_{5}(2)$, with $d=10$, $\ell=11$;
  7. (7)
    $3.M_{22}<{\operatorname{SU}}_{6}(2)$, with $d=10$, $\ell=11$; or
  8. (8)
    $3_{1}.{\operatorname{U}}_{4}(3).2_{2}<{\operatorname{SU}}_{6}(2)$, with $d=3$, $\ell=7$.
Proof.
First assume $S:=F^{*}(M)/{\mathbf{Z}}(F^{*}(M))$ is not of Lie type in the same characteristic as $G$. For $n\leq 12$ we again extract the relevant cases from the tables in [2, §8]. Note that $|{\operatorname{SU}}_{n}(q)|>|{\operatorname{SL}}_{n}(q)|$ for $n\geq 3$. Then arguing as in the proof of Lemma 4.3 we obtain that $|N|\geq q^{n-1}$ here as well. The possibilities for $S$ not excluded by this lower bound are now handled using the Atlas, leading to Cases (1)–(8). We may now assume $n\geq 13$, and so $N$ contains a torus of size at least $(q^{2}-1)^{6}\geq 3^{6}$ if $q\leq 3$, and $3^{11}$ if $q\geq 4$.
The considerations in the proof of Proposition 4.4 for $S$ alternating apply verbatim to show that no embeddings in dimension $n\geq 13$ lead to examples. In fact, the same is true for the whole discussion of cross characteristic embeddings of groups of Lie type as well as for embeddings of sporadic groups, again showing that no case with $n\geq 13$ arises.
If $S$ is of Lie type in defining characteristic, we use again [2, §8] when $n\leq 12$ to rule out the occurrence of an example. For $n\geq 13$ we can argue as in the case of ${\operatorname{SL}}_{n}(q)$ that the $\ell$-rank of any candidate $S$ is too small. ∎
Remark 4.9.
Again we see that in the situation of Proposition 4.7 the normaliser of a Sylow $d$-torus lies in a unique maximal subgroup of ${\operatorname{SU}}_{n}(q)$ whenever Sylow $\ell$-subgroups are non-cyclic, except in ${\operatorname{SU}}_{3}(5)$ with $\ell=3$ (where Sylow 3-subgroups are non-abelian). (Again, compare to [1].)
The $d$-split Levi subgroups of $G={\operatorname{SU}}_{n}(q)$ have the form
$\Big({\operatorname{GL}}_{n_{1}}((-q)^{e})\times\cdots\times{\operatorname{GL}}_{n_{t}}((-q)^{e})\times{\operatorname{GU}}_{s}(q)\Big)\cap G\quad\text{with $e\sum_{i}n_{i}+s=n$}$
by [9, Exmp. 3.5.14], where we may and will assume $n_{1}\geq\ldots\geq n_{t}$; note that here $s\equiv r\pmod{e}$ and thus $s\geq r$. For abbreviation we will write $L_{e}(n_{1},\ldots,n_{t};s)$ for such a Levi subgroup.
With this information in place we can determine the subnormalisers:
Theorem 4.10.
Let $G={\operatorname{SU}}_{n}(q)$ with $n\geq 3$, let $\ell\nmid q$ be a prime such that Sylow $\ell$-subgroups of $G$ are abelian and ${\mathbf{S}}_{d}\leq{\mathbf{G}}$ a Sylow $d$-torus, where $d:=e_{\ell}(q)$. With $e:=e_{\ell}(-q)$ write $n=ae+r$ with $0\leq r<e$. Then for $x\in{\mathbf{S}}_{d}^{F}$ an $\ell$-element we have one of
  1. (1)
    ${\operatorname{Sub}}_{G}(x)={\mathbf{N}}_{G}({\mathbf{S}}_{d})$ if ${\mathbf{C}}_{G}(x)={\mathbf{C}}_{G}({\mathbf{S}}_{d})$;
  2. (2)
    ${\operatorname{Sub}}_{G}(x)=({\operatorname{GU}}_{ae}(q)\times{\operatorname{GU}}_{r}(q))\cap G$ if $r>0$ and ${\mathbf{C}}_{G}(x)=L_{e}(n_{1},\ldots,n_{t};r)$ with $a=\sum n_{i}$ and at least one $n_{i}>1$;
  3. (3)
    ${\operatorname{Sub}}_{G}(x)=({\operatorname{GU}}_{e}(q)\wr{\mathfrak{S}}_{a})\cap G$ if $r=0$, $e>1$, $a>1$ and ${\mathbf{C}}_{G}(x)=L_{e}(1,\ldots,1;e)$; or
  4. (4)
    ${\operatorname{Sub}}_{G}(x)=G$ otherwise.
Proof.
We proceed as in the proof of Theorem 4.6. Observe that $\ell>2$ by our assumptions. By Proposition 2.2 the cyclic Sylow case again leads to (1). Now, for all maximal subgroups in class $\mathcal{S}$} coming up in Proposition 4.8, the Sylow $\ell$-subgroups of $G$ are either cyclic or non-abelian, so no further cases arise from these.
Now assume ${\operatorname{Sub}}_{G}(x)$ lies in one of the maximal subgroups $M$ listed in Proposition 4.7. Again by [5, Prop. 2.2], ${\mathbf{C}}_{G}(x)$ is a $d$-split Levi subgroup of $G$ and hence has the form $L_{e}(n_{1},\ldots,n_{t};s)$ introduce above, where either $s>r$ or $n_{1}>1$ as otherwise ${\mathbf{C}}_{G}(x)={\mathbf{C}}_{G}({\mathbf{S}}_{d})$ and we are in (1) of the conclusion. Now the groups in Cases (3), (4) and (5) in Proposition 4.7 have cyclic Sylow $\ell$-subgroup. If we are in Case (2), and so $r=0$, then arguing as for ${\operatorname{SL}}_{n}(q)$ we see that ${\operatorname{Sub}}_{G}(x)\leq M$ as in (3) of the conclusion if and only if ${\mathbf{C}}_{G}(x)$ is as claimed. Note here that when $e=1$ we obtain (1), and when $a=1$ we have ${\operatorname{Sub}}_{G}(x)=G$ as in (4). Finally, if $M$ is a Levi subgroup of $G$ as in Case (1) of Proposition 4.7 then necessarily $r=s$ for ${\mathbf{C}}_{G}(x)$ and ${\operatorname{Sub}}_{G}(x)=\langle{\mathbf{C}}_{G}(x),{\mathbf{N}}_{G}({\mathbf{S}}_{d})\rangle=M$ as in (2) of the conclusion by Proposition 2.1. ∎
4.3. The symplectic and odd-dimensional orthogonal groups
We now turn to $G$ either ${\operatorname{Sp}}_{2n}(q)$ or ${\operatorname{SO}}_{2n+1}(q)$ with $n\geq 2$, $(n,q)\neq(2,2)$. Let $d\geq 1$ be an integer and
$e:=\begin{cases}d&\text{if $d$ is odd},\\ d/2&\text{if $d$ is even}.\end{cases}$
We write $n=ae+r$ with $0\leq r<e$. Then the centraliser and normaliser of a Sylow $d$-torus ${\mathbf{S}}_{d}$ of the underlying simple algebraic group $\mathbf{G}$} are given as follows (see [9, Exmp. 3.5.15 and 3.5.29]):
${\mathbf{C}}_{G}({\mathbf{S}}_{d})={\operatorname{GL}}_{1}(q^{e})^{a}\times H_{r}\quad\text{and}\quad{\mathbf{N}}_{G}({\mathbf{S}}_{d})={\operatorname{GL}}_{1}(q^{e})\wr G(2e,1,a)\times H_{r}$
if $d=e$ is odd, and
${\mathbf{C}}_{G}({\mathbf{S}}_{d})={\operatorname{GU}}_{1}(q^{e})^{a}\times H_{r}\quad\text{and}\quad{\mathbf{N}}_{G}({\mathbf{S}}_{d})={\operatorname{GU}}_{1}(q^{e})\wr G(2e,1,a)\times H_{r}$
if $d=2e$ is even, where $H_{r}$ is a group of rank $r$ of the same classical type as $G$. We further assume there is a Zsigmondy prime divisor $\ell>2$ of $q^{d}-1$, so that $d=e_{\ell}(q)$, $e=e_{\ell}(q^{2})$.
Proposition 4.11.
Let $G={\operatorname{Sp}}_{2n}(q)$ with $n\geq 2$ and $d,e,\ell$ as above. If $M<G$ is a maximal subgroup containing the normaliser of a Sylow $d$-torus of $\mathbf{G}$}, then one of:
  1. (1)
    $M={\operatorname{Sp}}_{2ae}(q)\times{\operatorname{Sp}}_{2r}(q)$ if $r>0$;
  2. (2)
    $M={\operatorname{Sp}}_{2e}(q)\wr{\mathfrak{S}}_{a}$ if $n=ae$ and $a>1$;
  3. (3)
    $M={\operatorname{GL}}_{n}(q).2$ if $n=d$ is odd and $q$ is odd;
  4. (4)
    $M={\operatorname{Sp}}_{2n/t}(q^{t}).(q-1,2,t)$ if $n=e$ and $t|n$ is a prime;
  5. (5)
    $M={\operatorname{GU}}_{n}(q).2$ if $n=e=d/2$ is odd and $q$ is odd;
  6. (6)
    $M={\operatorname{Sp}}_{2n}(2)$ if $q=4$ and $d=1$;
  7. (7)
    $M={\operatorname{GO}}_{2n}^{+}(q)$ if $q$ is even, $n=ae>2$ and $2n/d$ is even;
  8. (8)
    $M={\operatorname{GO}}_{2n}^{-}(q)$ if $q$ is even, $n=ae>2$ and $2n/d$ is odd; or
  9. (9)
    $M$ is in class $\mathcal{S}$}.
Proof.
Let $N:={\mathbf{N}}_{G}({\mathbf{S}}_{d})$ for a Sylow $d$-torus ${\mathbf{S}}_{d}\leq{\mathbf{G}}$. By the description recalled above, $N$ acts irreducibly on subspaces of dimensions $ae,ae$ and $2r$ of the natural module of $G$. We refer to [12, Tab. 3.5.C] for the Aschbacher classes.
The maximal parabolic subgroups $P_{r}$ (for $r>0$) do not contain a Sylow $d$-torus of $G$ centralised by an ${\operatorname{Sp}}_{2r}(q)$-factor. Thus among reducible maximal subgroups we only obtain the one in (1). For the maximal imprimitive subgroups ${\operatorname{Sp}}_{2n/t}(q)\wr{\mathfrak{S}}_{t}$ with $2\leq t|n$, looking at the possible imprimitivity decompositions for $N$ we arrive at the case in (2). The imprimitive subgroup $M={\operatorname{GL}}_{n}(q).2$ with $q$ odd contains an elementary abelian $\ell$-subgroup of the required rank only when $d$ is odd. Its automiser is $G(2d,1,a)$ in $G$, and $G(d,1,a).2$ in $M$ by the description in Section 4.1, which agree only when $a=1$. Comparing the centralisers in $M$ and $G$ we see that then necessarily $d=n$, as in (3) of our conclusion.
The extension field subgroups $M={\operatorname{Sp}}_{2m}(q^{t}).(q-1,2,t)$, with $n=mt$ and $t$ prime contain an elementary abelian $\ell$-subgroup $E$ of the right rank only when $t|d$, respectively if $4|d$ when $t=2$. Now the automiser of $E$ in $M$ is G(2e/t,1,a).(q1,2,t)G(2e/t,1,a).(q-1,2,t) and $G(2e,1,a)$ in $G$, so we need $a=1$. Moreover, comparing centralisers we see $r=0$, whence $n=e$ as in (4). A subgroup $M={\operatorname{GU}}_{n}(q).2$, with $q$ odd, contains $E$ only when $d\equiv 2\pmod{4}$, so $e=d/2$ is odd, by the order formula. The automiser of $E$ in $M$ is now $G(e,1,a).2$, and $G(d,1,a)$ in $G$, forcing $a=1$. Comparison of centralisers shows $r=0$ and so $n=e$, as in (5). For the subfield subgroups ${\operatorname{Sp}}_{2n}(q_{0})$ with $q=q_{0}^{t}$, $t$ prime, we can argue as for ${\operatorname{SL}}_{n}(q)$ to arrive at $q=4$ and $d=1$, which in this case does lead to a containment, listed in (6).
The tensor product stabilisers of type ${\operatorname{Sp}}_{2m}(q){\operatorname{GO}}_{n/m}^{(\pm)}(q)$, with $m|n$, $n/m\geq 3$ and $q$ odd, have strictly smaller $\ell$-rank than $G$ for all relevant primes $\ell$. For $M$ the normalizer of an extra-special subgroup of type $2^{1+2m}.{\operatorname{GO}}_{2m}^{-}(2)$, with $n=2^{m-1}$ and $q=p$ odd, the size of an abelian subgroup of $M$ containing our elementary abelian $\ell$-subgroup is at most $3\cdot 2^{m(m-1)/2+m+1}$ by [25, Thm 3.1], while it is at least $(q-1)^{n}$ in $G$, respectively $7^{n/2}$ if $q=3$. This forces $m=4$ and $q=3$, so $M=2^{1+8}.{\operatorname{GO}}_{8}^{-}(2)$ in $G={\operatorname{Sp}}_{16}(3)$. Here, only $\ell=17$ is possible, but elements of order 17 in $G$ have too large a centraliser, so no example arises here.
Comparing orders it is easily seen that the tensor induced subgroups ${\operatorname{Sp}}_{m}(q)\wr{\mathfrak{S}}_{t}$ with $2n=m^{t}$, $t\geq 3$ odd, cannot contain an elementary abelian $\ell$-subgroup of the required rank. Finally, for $M={\operatorname{GO}}_{2n}^{\pm}(q)$ with $q$ odd, comparing centralisers of elementary abelian subgroups in Sylow $d$-tori, we arrive at the stated conditions in (7) and (8). In these cases, $M$ does indeed contain a conjugate of $N$ as can be seen from the description of the Sylow $d$-normalisers in orthogonal groups recalled in Section 4.4 below. ∎
We need not discuss the orthogonal groups ${\operatorname{SO}}_{2n+1}(q)$ for even $q$, since these are isomorphic to ${\operatorname{Sp}}_{2n}(q)$, nor ${\operatorname{SO}}_{5}(q)$ which possesses the same non-abelian simple composition factor as ${\operatorname{Sp}}_{4}(q)$ and otherwise only composition factors of order 2, whence subnormalisers of $\ell$-elements, for $\ell\neq 2$, of the two groups determine each other by [16, Lemma 2.15].
Proposition 4.12.
Let $G={\operatorname{SO}}_{2n+1}(q)$ with $n\geq 3$ and $q$ odd, and $d,e,\ell$ as above. If $M<G$ is a maximal subgroup containing the normaliser of a Sylow $d$-torus of $\mathbf{G}$}, then one of:
  1. (1)
    $M=({\operatorname{GO}}_{2ae}^{+}(q)\times{\operatorname{GO}}_{2r+1}(q))\cap G$ if $2ae/d$ is even;
  2. (2)
    $M=({\operatorname{GO}}_{2ae}^{-}(q)\times{\operatorname{GO}}_{2r+1}(q))\cap G$ if $2ae/d$ is odd; or
  3. (3)
    $M$ is in class $\mathcal{S}$}.
Proof.
Let $N:={\mathbf{N}}_{G}({\mathbf{S}}_{d})$ for a Sylow $d$-torus ${\mathbf{S}}_{d}\leq{\mathbf{G}}$. Here, by the structure recalled above, $N$ has irreducible constituents of dimensions $ae,ae$ and $2r+1$ on the natural module. We refer to [12, Tab. 3.5.D] for the Aschbacher classes.
The maximal parabolic subgroups $P_{r}$ (for $r>0$) do not possess an ${\operatorname{SO}}_{2r+1}(q)$-factor centralising a Sylow $d$-torus of $G$, so the only reducible cases are those in (1) and (2). By the order formulas, none of the other types of geometric subgroups can contain our subgroup $N$. For this, observe that all but the subfield subgroups are of strictly smaller rank than $G$. ∎
Lemma 4.13.
Let $n\geq 2$ and $q$ be a prime power. Then
$|{\operatorname{Sp}}_{2n}(q)|=|{\operatorname{SO}}_{2n+1}(q)|>(q^{2}-1)^{2}q^{2n^{2}+n-4}.$
Proof.
The order formula shows that $|{\operatorname{Sp}}_{2n}(q)|_{q^{\prime}}=(q^{2}-1)|{\operatorname{SL}}_{n}(q^{2})|_{q^{\prime}}$, thus an application of Lemma 4.2 gives the claim. ∎
Lemma 4.14.
For $G={\operatorname{Sp}}_{2n}(q)$ or ${\operatorname{SO}}_{2n+1}(q)$ with $n\geq 2$ the subgroup $N$ has order $|N|\geq 2^{a}q^{n}$, where $n=ae+r$ with $0\leq r<e$.
Proof.
Arguing as in the proof of Lemma 4.3 we see
$(q^{e}\pm 1)^{a}|G(2e,1,a)|\geq((3/4)q)^{e}(2e)^{a}a!\geq 2^{a}q^{ae}$
and then Lemma 4.13 yields the stated lower bound. ∎
We phrase the next result for the derived subgroup $\Omega_{2n+1}(q)=[{\operatorname{SO}}_{2n+1}(q),{\operatorname{SO}}_{2n+1}(q)]$ of index 2 in ${\operatorname{SO}}_{2n+1}(q)$ since there the normalisers of simple subgroups are more easily understood. As all Sylow $d$-tori are conjugate, $|{\mathbf{N}}_{{\operatorname{SO}}_{2n+1}(q)}({\mathbf{S}}_{d}):{\mathbf{N}}_{\Omega_{2n+1}(q)}({\mathbf{S}}_{d})|$ as well.
Proposition 4.15.
Let $G={\operatorname{Sp}}_{2n}(q)$ with $n\geq 2$, or $G=\Omega_{2n+1}(q)$ with $n\geq 3$ and $q$ odd and assume $M\in{\mathcal{S}}$ contains the normaliser $N$ of a Sylow $d$-torus. Then $M<G$ is one of:
  1. (1)
    ${\operatorname{U}}_{3}(3).2<{\operatorname{Sp}}_{6}(2)$, with $d=3$, $\ell=7$;
  2. (2)
    $2.{\operatorname{L}}_{2}(13)<{\operatorname{Sp}}_{6}(3)$, with $d=3$, $\ell=13$;
  3. (3)
    ${\operatorname{L}}_{2}(17)<{\operatorname{Sp}}_{8}(2)$, with $d=8$, $\ell=17$;
  4. (4)
    ${\mathfrak{S}}_{10}<{\operatorname{Sp}}_{8}(2)$, with $d=4$, $\ell=5$, or $d=3$, $\ell=7$;
  5. (5)
    ${\mathfrak{S}}_{14}<{\operatorname{Sp}}_{12}(2)$, with $d=3$, $\ell=7$;
  6. (6)
    ${\mathfrak{S}}_{9}<\Omega_{7}(3)$, with $d=4$, $\ell=5$, or $d=6$, $\ell=7$; or
  7. (7)
    $G_{2}(3)<\Omega_{7}(3)$, with $d=3$, $\ell=13$.
Proof.
Let $S:=F^{*}(M)/{\mathbf{Z}}(F^{*}(M))$. Again we start with the case that $S$ is not of Lie type in characteristic dividing $q$. As before, the case $n\leq 6$ for the symplectic groups respectively $n\leq 5$ for the orthogonal groups can be discussed using the tables in [2, §8], which leads to the groups in (1)–(6) of our conclusion. So we will assume now that $n\geq 7$ (resp. $n\geq 6$) and thus $N$ contains a maximal (abelian) torus of order at least $(q-1)^{6}\geq 3^{6}$ when $q\geq 4$, at least $(q^{2}-q+1)^{3}=7^{3}$ when $q=3$ and at least $(q^{2}-1)^{4}=3^{4}$ when $q=2$ (using that $q$ is odd when $G=\Omega_{2n+1}(q)$).
First assume $S={\mathfrak{A}}_{m}$. Since $2n+1\geq 13$ we then have $m\leq 2n+2$. The normaliser $N$ contains a maximal torus $T$ of order at least $(q-1)^{n}$, so at least $3^{n}$ if $q\geq 4$. Since maximal abelian subgroups of ${\mathfrak{S}}_{m}$ have size at most $3^{m/3}\leq 3^{(2n+2)/3}$, comparison shows we must have $q\leq 3$. If $q=3$ then $|T|=(q^{e}\pm 1)^{a}(q+1)^{r}\geq 8^{n/2}$ and comparing with the bound for ${\mathfrak{S}}_{m}$ we arrive at $n\leq 2.4<6$. Assume $q=2$, and so $G$ is symplectic. If $\ell=3$, so $d=2$, then $|T|=3^{n}$ which is always larger than $3^{(2n+2)/3}$. When $\ell\geq 5$ observe that $\ell$-elements in $T$ are conjugate to $2e$ of their powers, while $\ell$-elements in ${\mathfrak{S}}_{m}$ are rational, so we conclude that $\ell=2e+1$. Now a maximal elementary abelian $\ell$-subgroup $E$ of ${\mathfrak{S}}_{m}$ has rank $\lfloor m/\ell\rfloor$ and the largest abelian subgroup containing it has order at most $3^{m^{\prime}}|E|$ with $m^{\prime}=(m-\ell\lfloor m/\ell\rfloor)/3$, while in $G$ such a subgroup has rank $a=\lfloor n/e\rfloor$ and is contained in a torus of order at least $(2^{e}-1)^{a}(2^{r}+1)$. Comparing the two we conclude that $\ell=5$ and $7\leq n\leq 9$. In all of these cases, the centraliser of $E$ in ${\mathfrak{S}}_{m}$ does not contain a subgroup ${\operatorname{Sp}}_{2r}(2)$.
If $S$ is sporadic, since $N$ contains a cyclic subgroup of order $q^{e}\pm 1$, as in the proof of Proposition 4.4 we conclude that $e\leq 6$, and in fact $e\leq 4$ when $q\geq 3$, and $e\leq 3$ when $q\geq 4$. Since no covering group of a sporadic simple group possesses elements of order 65, 80 or 82, we have $e\leq 5$ when $q=2$ and $e\leq 3$ when $q\geq 4$. Now when $e\leq 3$ then $a\geq 2$ as $n\geq 7$ and hence $\ell^{2}$ divides $|S|$. In this case, as well as when $q=2$, $e=4,5$ and $7\leq n\leq 9$, the table of low dimensional representations in [11] shows that no example arises.
Next assume $S$ if of Lie type in cross characteristic. For $S={\operatorname{L}}_{m}(y)$ with $m\geq 3$, as in the proof of Proposition 4.4 the smallest degree is $(y^{m}-1)/(y-1)-m\geq y^{m-1}$, unless $(m,y)\in{\mathcal{E}}:=\{(3,2),(3,4),(4,2),(4,3)\}$, and we have $|M|\leq|{\operatorname{Aut}}(S)|\leq 2|{\operatorname{PGL}}_{m}(y)|\log y\leq 2\ y^{m^{2}-1}\log y$. Thus, for $(m,y)\notin{\mathcal{E}}$, if $N\leq M$ then
$2q^{y^{m-1}/2}\leq 2^{a}q^{n}\leq|N|\leq|M|\leq 2y^{m^{2}-1}\log y$
by Lemma 4.14. Assuming $2n+1\geq 13$, this is only satisfied for $S={\operatorname{L}}_{3}(5),{\operatorname{L}}_{5}(2),{\operatorname{L}}_{6}(2)$, with $2n+1$ at most 19, 14, 21. But none of these groups has a faithful projective representation in this range, by [11]. The groups with $(m,y)\in{\mathcal{E}}$ can be excluded in a similar way, so no further case arises.
For $S={\operatorname{L}}_{2}(y)$, $y\geq 7$ and $y\neq 9$, the Sylow $\ell$-subgroups of $S$ are cyclic, with centraliser order at most $y+1$, while the centraliser of an $\ell$-element in $T\leq G$ has size at least $q^{(n+1)/2}-1$, so $y\geq q^{(n+1)/2}-2$. On the other hand, the minimal projective degree of $S$ is (y1)/(2,y1)(y-1)/(2,y-1), so we need $n\geq(y-1)/2$. This is only satisfied for values $n\leq 7$ which does not lead to new cases. The other series of groups can be dealt with by analogues estimates.
If $S$ is of Lie type in characteristic $p$, the low dimensional cases can again be discussed using [2, §8], which leads only to the example in (7) of the conclusion. So assume $n\geq 7$, respectively $n\geq 6$ for $G=\Omega_{2n+1}(q)$. If $S={\operatorname{L}}_{m}(p^{f})$ then by [12, Prop. 5.4.6, 5.4.8] we must have $2n+1\geq m(m-1)$ (since the representation needs to be self-dual) and $p^{f}|q$, or $2n+1\geq m^{k}$ and $p^{f}|q^{k}$ for some $k\geq 2$. Write $m=bd+s$, so $S$ has $\ell$-rank at most $b$, then $2n+1\geq(bd+s)(bd+s-1)$ and since $d\geq e$ we find $n\geq e(eb^{2}/2+bs-b/2)$ whence $G$ has $\ell$-rank at least $eb^{2}/2+bs-b/2$. This is larger than $b$ unless $m\leq 6$, and in those cases we obtain $n\leq 5$, which was excluded. Again the same type of estimates applies to all other types of $S$, leading to no further candidates. ∎
Remark 4.16.
(1) The maximal subgroups of $\Omega_{7}(3)$ in (6) and (7) of the preceding result are not stable under the diagonal automorphism induced by ${\operatorname{SO}}_{7}(3)$ (see [2, Tab. 8.40]).
(2) The above shows that for the symplectic groups ${\operatorname{Sp}}_{2n}(q)$ with $q$ even it may happen that the normaliser of a Sylow $d$-torus lies in two distinct maximal subgroups even when Sylow $\ell$-subgroups are non-cyclic, namely in Cases (2) and (7), (8) of Proposition 4.11. Further such cases occur for $\ell\leq 7$. For ${\operatorname{SO}}_{2n+1}(q)$ with $q$ odd, there is always a unique maximal overgroup in the non-cyclic Sylow case.
We next determine the subnormalisers. For this we recall from [9, Exmp. 3.5.15] that the $d$-split Levi subgroups of $G$ have the form
${\operatorname{GL}}_{n_{1}}(q^{e})\times\cdots\times{\operatorname{GL}}_{n_{t}}(q^{e})\times H_{s}\qquad\text{if $d=e$ is odd},$
respectively
${\operatorname{GU}}_{n_{1}}(q^{e})\times\cdots\times{\operatorname{GU}}_{n_{t}}(q^{e})\times H_{s}\qquad\text{if $d=2e$ is even},\quad$
where $H_{s}={\operatorname{Sp}}_{2s}(q)$ resp. ${\operatorname{SO}}_{2s+1}(q)$ and $e\sum_{i}n_{i}+s=n$ in either case, where we may assume $n_{1}\geq n_{2}\geq\cdots\geq n_{t}$. Here again $s\equiv r\pmod{e}$ and thus $s\geq r$. We will write $L_{d}(n_{1},\ldots,n_{t};s)$ for such a Levi subgroup (whose structure depends on the parity of $d$ as indicated).
Theorem 4.17.
Let $G={\operatorname{Sp}}_{2n}(q)$ with $n\geq 2$, let $\ell\nmid q$ be a prime such that Sylow $\ell$-subgroups of $G$ are abelian and ${\mathbf{S}}_{d}\leq{\mathbf{G}}$ a Sylow $d$-torus, where $d:=e_{\ell}(q)$. With $e:=e_{\ell}(q^{2})$ write $n=ae+r$ where $0\leq r<e$. Then for $x\in{\mathbf{S}}_{d}^{F}$ an $\ell$-element we have one of
  1. (1)
    ${\operatorname{Sub}}_{G}(x)={\mathbf{N}}_{G}({\mathbf{S}}_{d})$ if ${\mathbf{C}}_{G}(x)={\mathbf{C}}_{G}({\mathbf{S}}_{d})$;
  2. (2)
    ${\operatorname{Sub}}_{G}(x)={\operatorname{Sp}}_{2ae}(q)\times{\operatorname{Sp}}_{2r}(q)$ if $r>0$, $q$ is odd and ${\mathbf{C}}_{G}(x)=L_{d}(n_{1},\ldots,n_{t};r)$ with $n_{1}>1$;
  3. (3)
    ${\operatorname{Sub}}_{G}(x)={\operatorname{GO}}_{2ae}^{+}(q)\times{\operatorname{Sp}}_{2r}(q)$ if $q$ is even, $2ae/d$ is even and ${\mathbf{C}}_{G}(x)=L_{d}(n_{1},\ldots,n_{t};r)$ with $n_{1}>1$;
  4. (4)
    ${\operatorname{Sub}}_{G}(x)={\operatorname{GO}}_{2ae}^{-}(q)\times{\operatorname{Sp}}_{2r}(q)$ if $q$ is even, $2ae/d$ is odd and ${\mathbf{C}}_{G}(x)=L_{d}(n_{1},\ldots,n_{t};r)$ with $n_{1}>1$;
  5. (5)
    ${\operatorname{Sub}}_{G}(x)={\operatorname{Sp}}_{2e}(q)\wr{\mathfrak{S}}_{a}$ if $r=0$, $a>1$ and ${\mathbf{C}}_{G}(x)=L_{d}(1,\ldots,1;e)$; or
  6. (6)
    ${\operatorname{Sub}}_{G}(x)=G$ otherwise.
Proof.
The cyclic Sylow case is again covered by Proposition 2.2. So assume Sylow $\ell$-subgroups of $G$ are non-cyclic and ${\operatorname{Sub}}_{G}(x)$ is contained in a maximal subgroup $M$ of $G$. Among the maximal subgroups in Proposition 4.15 only ${\mathfrak{S}}_{10}<{\operatorname{Sp}}_{8}(2)$ with $\ell=5$ and ${\mathfrak{S}}_{14}<{\operatorname{Sp}}_{12}(2)$ with $\ell=7$ possess non-cyclic Sylow $\ell$-subgroups. Direct computation in GAP shows that in either case, the subnormalisers of $\ell$-elements are as in (3) or (5) of the conclusion (and hence not symmetric groups).
Now assume ${\mathbf{C}}_{G}(x)$, and hence ${\operatorname{Sub}}_{G}(x)$, is contained in one of the maximal subgroups $M$ in Proposition 4.11. Again by [5, Prop. 2.2] the centraliser ${\mathbf{C}}_{\mathbf{G}}(x)$ is a $d$-split Levi subgroup of $\mathbf{G}$}. If $r=0$ then Cases (2)–(8) in Proposition 4.11 are relevant. In Cases (3), (4) and (5) the Sylow $\ell$-subgroups of $G$ are cyclic. In Case (6) we have $\ell=3$ and Sylow $3$-subgroups are abelian only when $n=2$, a possibility that can be ruled out by explicit computation. If ${\mathbf{C}}_{G}(x)$ lies in $M={\operatorname{Sp}}_{2e}(q)\wr{\mathfrak{S}}_{a}$ then as in the linear and unitary cases necessarily $s=e$, while all $n_{i}=1$, giving (5). If ${\mathbf{C}}_{G}(x)$ lies in an orthogonal group as in Cases (7) or (8) then comparing centralisers shows that $s=0$, leading to the conclusion in (3) or (4).
Finally, if $r>0$ then the only maximal overgroup for $N$ is a subsystem subgroup $M={\operatorname{Sp}}_{2ae}(q)\times{\operatorname{Sp}}_{2r}(q)$ as in Case (1) of Proposition 4.11. Comparing dimensions of composition factors on the natural module we conclude that $r=s$, and in this case, since $M$ contains the normaliser of a Sylow $d$-torus, ${\operatorname{Sub}}_{G}(x)\leq M$ by Proposition 2.1. Now going again through the cases in Proposition 4.11 for the ${\operatorname{Sp}}_{2ae}(q)$-factor it follows that we must in fact have ${\operatorname{Sub}}_{G}(x)=M$ when $q$ is odd, while when $q$ is even, ${\mathbf{C}}_{G}(x)$ and hence ${\operatorname{Sub}}_{G}(x)$ lies in a subgroup as given in (3) or (4). ∎
Theorem 4.18.
Let $G={\operatorname{SO}}_{2n+1}(q)$ with $n\geq 3$ and $q$ odd, let $\ell\nmid q$ be a prime such that Sylow $\ell$-subgroups of $G$ are abelian and ${\mathbf{S}}_{d}\leq{\mathbf{G}}$ a Sylow $d$-torus, where $d:=e_{\ell}(q)$. With $e:=e_{\ell}(q^{2})$ write $n=ae+r$ where $0\leq r<e$. Then for $x\in{\mathbf{S}}_{d}^{F}$ an $\ell$-element one of
  1. (1)
    ${\operatorname{Sub}}_{G}(x)={\mathbf{N}}_{G}({\mathbf{S}}_{d})$ if ${\mathbf{C}}_{G}(x)={\mathbf{C}}_{G}({\mathbf{S}}_{d})$;
  2. (2)
    ${\operatorname{Sub}}_{G}(x)=\big({\operatorname{GO}}_{2ae}^{+}(q)\times{\operatorname{GO}}_{2r+1}(q)\big)\cap G$ if $2ae/d$ is even and ${\mathbf{C}}_{G}(x)=L_{d}(n_{1},\ldots,n_{t};r)$ with $n_{1}>1$;
  3. (3)
    ${\operatorname{Sub}}_{G}(x)=\big({\operatorname{GO}}_{2ae}^{-}(q)\times{\operatorname{GO}}_{2r+1}(q)\big)\cap G$ if $2ae/d$ is odd and ${\mathbf{C}}_{G}(x)=L_{d}(n_{1},\ldots,n_{t};r)$ with $n_{1}>1$; or
  4. (4)
    ${\operatorname{Sub}}_{G}(x)=G$ otherwise.
Proof.
The proof is very similar to the one of Proposition 4.17, but easier, as there are fewer cases to consider: The two groups in Proposition 4.15(6) and (7) possess cyclic Sylow $\ell$-subgroups and thus do not occur as subnormalisers by Proposition 2.2 while the maximal subgroups in (1) and (2) of Proposition 4.12 lead to (2) and (3) of the conclusion. ∎
4.4. The even-dimensional orthogonal groups
Finally, we consider the even-dimensional special orthogonal groups ${\operatorname{SO}}_{2n}^{\pm}(q)$, $n\geq 4$. Here, by convention we let ${\operatorname{SO}}_{2n}:={\operatorname{GO}}_{2n}^{\circ}$, the connected component of the identity, and ${\operatorname{SO}}_{2n}^{\pm}(q)$ the group of fixed points under a Frobenius map $F$ with respect to an ${\mathbb{F}}_{q}$-structure.
Let $d\geq 1$ be an integer and as before set $e:=d$ if $d$ is odd, $e:=d/2$ if $d$ is even. We write $n=ae+r$ with $0\leq r<e$, except if $d|2n$ and either $2n/d$ is odd in type ${\operatorname{SO}}_{2n}^{+}(q)$, or $2n/d$ is even in type ${\operatorname{SO}}_{2n}^{-}(q)$: in the latter cases, write $n=(a+1)e$ and set $r:=e$ (note that $e|n$ under our assumptions). Then the centraliser and normaliser of a Sylow $d$-torus ${\mathbf{S}}_{d}$ of ${\mathbf{G}}={\operatorname{SO}}_{2n}$ in $\hat{G}:={\operatorname{GO}}_{2n}^{\pm}(q)$ are given as follows (see [9, Exmp. 3.5.15 and 3.5.29]):
${\mathbf{C}}_{\hat{G}}({\mathbf{S}}_{d})={\operatorname{GL}}_{1}(q^{e})^{a}\times H_{r}\quad\text{and}\quad{\mathbf{N}}_{\hat{G}}({\mathbf{S}}_{d})={\operatorname{GL}}_{1}(q^{e})\wr G(2e,1,a)\times H_{r}$
if $d=e$ is odd, and
${\mathbf{C}}_{\hat{G}}({\mathbf{S}}_{d})={\operatorname{GU}}_{1}(q^{e})^{a}\times H_{r}\quad\text{and}\quad{\mathbf{N}}_{\hat{G}}({\mathbf{S}}_{d})={\operatorname{GU}}_{1}(q^{e})\wr G(2e,1,a)\times H_{r}$
if $d=2e$ is even, where $H_{r}$ is a general orthogonal group of rank $r$ of the same type as $G$, except that it is of opposite sign if $d$ is even and $a$ is odd. Again, we assume there is a Zsigmondy prime divisor $\ell>2$ of $q^{d}-1$, so that $d=e_{\ell}(q)$, $e=e_{\ell}(q^{2})$.
We again first consider the maximal overgroups of Sylow tori normalisers:
Proposition 4.19.
Let $G={\operatorname{SO}}_{2n}^{+}(q)$ with $n\geq 4$, and $d,e,\ell$ as above. If $M<G$ is a maximal subgroup containing the normaliser of a Sylow $d$-torus, then one of:
  1. (1)
    $M=({\operatorname{GO}}_{2ae}^{+}(q)\times{\operatorname{GO}}_{2r}^{+}(q))\cap G$ if $r>0$, and $a$ is even or $d$ is odd;
  2. (2)
    $M=({\operatorname{GO}}_{2ae}^{-}(q)\times{\operatorname{GO}}_{2r}^{-}(q))\cap G$ if $r>0$, $a$ is odd and $d$ is even;
  3. (3)
    $M={\operatorname{Sp}}_{2n-2}(2)$ if $q=2$, $r=1$, and $a$ is even or $d>1$ is odd;
  4. (4)
    $M=({\operatorname{GO}}_{2e}^{+}(q)\wr{\mathfrak{S}}_{a})\cap G$ if $r=0$, $a>1$ and $d$ is odd;
  5. (5)
    $M=({\operatorname{GO}}_{2e}^{-}(q)\wr{\mathfrak{S}}_{a})\cap G$ if $r=0$, $a$ is even and $d$ is even;
  6. (6)
    $M={\operatorname{GL}}_{n}(q).2$ if $e=d$ is odd, and $n=e+1$ or $n=2e$;
  7. (7)
    $M={\operatorname{GU}}_{n}(q).2$ if $e=d/2$ is odd, and $n=e+1$ or $n=2e$;
  8. (8)
    $M={\operatorname{GO}}_{2n/t}^{+}(q^{t}).t\cap G$ if $n=d$ is odd, $2<t|n$ is prime and $2n/t\geq 4$;
  9. (9)
    $M={\operatorname{GO}}_{n}^{+}(q^{2}).2\cap G$ if $n=d\equiv 0\pmod{4}$;
  10. (10)
    $M={\operatorname{GO}}_{2n}^{+}(2)$ if $q=4$, $d=1$ and $n$ is even;
  11. (11)
    $M={\operatorname{GO}}_{2n}^{-}(2)$ if $q=4$, $d=1$ and $n$ is odd; or
  12. (12)
    $M$ is in class $\mathcal{S}$}.
Observe that in items (4) and (5) for $e=1$, so $d\in\{1,2\}$, we have $M=N$, hence $N$ is maximal in those cases.
Proof.
Let ${\mathbf{S}}_{d}$ be a Sylow $d$-torus of $({\mathbf{G}},F)$. The structure of ${\mathbf{C}}_{G}({\mathbf{S}}_{d})$ and $N:={\mathbf{N}}_{G}({\mathbf{S}}_{d})$ was recalled above. We refer to [12, Tab. 3.5.E] and the relevant tables in [2] for a description of the Aschbacher classes. The composition factors of $N$ on the natural module have dimensions $ae,ae$ and $2r$. From this it follows that reducible maximal subgroups containing $N$ are either as in (1), (2), or, if $N$ lies in a subgroup $M={\operatorname{Sp}}_{2n-2}(q)$ with $q$ even, then we must have 2n2=2ae2n-2=2ae and thus $r=1$. In this case, the centraliser of ${\mathbf{S}}_{d}^{F}$ in $G$ is larger than the one in $M$ by a factor ${\operatorname{SO}}_{2}^{\epsilon}(q)$, which forces $q=2$ and $\epsilon=+$, and the latter holds when $a$ is even or $d$ is odd, giving (3).
The maximal imprimitive subgroups lead to cases (4)–(6). Here note that normalisers of subgroups of type ${\operatorname{GL}}_{n}(q)$ are only maximal in ${\operatorname{SO}}_{2n}^{+}(q)$ when $n$ is even. In this case, they contain a Sylow $d$-torus of $G$ when $d\geq 3$ is odd, and its full normaliser only when moreover either $n=2d$ or $n=d+1$. Further, the imprimitive maximal subgroup ${\operatorname{GO}}_{n}(q)\wr 2$, with $nq$ odd, is of smaller rank than $G$ and thus the centraliser of its Sylow $d$-torus is smaller than the one in $G$ (using the precise descriptions in [12, Prop. 4.2.14, 4.2.16]). The same holds for the extension field subgroup ${\operatorname{GO}}_{n}(q^{2})$, $nq$ odd, using [12, Prop. 4.3.20].
Normalisers of subgroups of type ${\operatorname{GU}}_{n}(q)$, with $n$ even, contain a Sylow $d$-torus only when $d\equiv 2\pmod{4}$, and the full normaliser of ${\mathbf{S}}_{d}$ only when either $n=d$ or $n=d/2+1$, giving (7). For $M$ of type ${\operatorname{GO}}_{2n/t}^{+}(q^{t})$, with $t|n$ a prime and $n/t\geq 2$, to contain a Sylow $d$-torus, we need $t|d$, and $4|d$ if $t=2$. Comparing centralisers yields that $d$ must divide $n$. Now if $t$ is odd, the automiser of ${\mathbf{S}}_{d}$ in $M$ is $G(2e/t,2,a).t$ and $G(2e,2,a)$ in $G$, forcing $a=1$, so $n=e$ must be odd (as $d|n$). If $t=2$ then $4|d|n$ forces $a\geq 2$. The automiser of ${\mathbf{S}}_{d}$ in $M$ is $G(d/2,1,a).2$ but $G(d,2,a)$ in $G$, which agree only when $a=2$, so $n=2e=d$. We hence reach the conditions stated in (8) and (9).
For the subfield subgroups, comparing orders of maximal tori as in the earlier proofs, we see that necessarily $q=4$ and $d=1$, and thus only the groups in (10) or (11) can contain the normaliser of a Sylow $d$-torus. If $M$ is the normaliser of an extra-special 2-group, with $q$ odd, estimates as in the proof for ${\operatorname{Sp}}_{2n}(q)$ rule out all cases with $m\geq 4$ except $M=2^{1+2m}.{\operatorname{GO}}_{2m}^{+}(2)<{\operatorname{SO}}_{2^{m}}^{+}(3)$ for $m=4,5$. For those, the only prime $\ell$ such that $M$ contains a Sylow $\ell$-subgroup of $G$ is $\ell=31$ when $m=5$, but then its centraliser in $G$ is much too big.
Finally, as in the other types, the tensor product subgroups and the tensor induced subgroups have too small a rank to contain an $\ell$-group of the necessary rank. ∎
Proposition 4.20.
Let $G={\operatorname{SO}}_{2n}^{-}(q)$ with $n\geq 4$ and let $d,e,\ell$ be as above. If $M<G$ is a maximal subgroup containing the normaliser of a Sylow $d$-torus, then one of:
  1. (1)
    $M=({\operatorname{GO}}_{2ae}^{+}(q)\times{\operatorname{GO}}_{2r}^{-}(q))\cap G$ if $r>0$, and $a$ is even or $d$ is odd;
  2. (2)
    $M=({\operatorname{GO}}_{2ae}^{-}(q)\times{\operatorname{GO}}_{2r}^{+}(q))\cap G$ if $r>0$, $a$ is odd and $d$ is even;
  3. (3)
    $M={\operatorname{Sp}}_{2n-2}(2)$ if $q=2$, $r=1$, $a$ is odd and $d$ is even;
  4. (4)
    $M=({\operatorname{GO}}_{2e}^{-}(q)\wr{\mathfrak{S}}_{a})\cap G$ if $r=0$, $a$ is odd and $d$ is even;
  5. (5)
    $M={\operatorname{GU}}_{n}(q)$ if $n=e=d/2$ is odd;
  6. (6)
    $M={\operatorname{GO}}_{n}(9)$ if $q=3$, $d\equiv 0\pmod{4}$ and $n=d/2+1$;
  7. (7)
    $M={\operatorname{GO}}_{2n/t}^{-}(q^{t}).t\cap G$ if $n=e=d/2$, $t|n$ is prime and $2n/t\geq 4$ is even; or
  8. (8)
    $M$ is in class $\mathcal{S}$}.
Proof.
We use the description in [12, Tab. 3.5.F] of the Aschbacher classes. Among reducible maximal subgroups we only find the examples listed in (1)–(3), with arguments entirely similar to the ones used for the groups of plus-type. Next, using the dimensions of the $N$-composition factors on the natural module, we find the imprimitive subgroups in (4). Here, the subgroup ${\operatorname{GO}}_{n}(q)^{2}$, with $nq$ odd, is of smaller rank than $G$ and using the description in [12, Prop. 4.1.6,4.2.16] cannot contain a Sylow $d$-torus normaliser.
The subgroup $M={\operatorname{GU}}_{n}(q)$ with $n$ odd contains a suitable elementary abelian $\ell$-subgroup only when $d\equiv 2\pmod{4}$ and thus $e=d/2$ is odd. Its automiser in $M$ is then $G(e,1,a)$, and in $G$ it equals $G(2e,2,a)$, so we must have $a=1$. Comparing centralisers shows $r=0$, hence $n=e=d/2$ as listed in (5). The extension field subgroups ${\operatorname{GO}}_{n}(q^{2})$ with $nq$ odd do only occur if $q=3$ and d=2n2d=2n-2, as listed in (6), taking into account the information in [12, Prop. 4.3.20], while the extension field subgroups ${\operatorname{GO}}_{2n/t}^{-}(q^{t}).t$ with $t|2n$ prime, $2n/t\geq 3$, can be discussed as in the earlier proofs, leading to (7) of the conclusion. The tensor product subgroups have smaller rank than necessary, and the maximal tori in subfield subgroups of type ${\operatorname{GO}}_{2n}^{-}(q_{0})$, with $q=q_{0}^{t}$ and $t\geq 3$ prime, are too small. ∎
In order to deal with the maximal subgroups in class $\mathcal{S}$} we again first derive a lower bound for $|N|$:
Lemma 4.21.
Let $n\geq 4$ and $q$ be a prime power. Then
$|{\operatorname{SO}}_{2n}^{\pm}(q)|>\frac{1}{2}q^{2n^{2}-n}.$
Proof.
Since $|{\operatorname{SO}}_{2n}^{-}(q)|>|{\operatorname{SO}}_{2n}^{+}(q)|$ it suffices to consider the latter group. By the order formula $|{\operatorname{SO}}_{2n}^{+}(q)|=q^{n-1}(q^{n}-1)|{\operatorname{Sp}}_{2n-2}(q)|$ so from Lemma 4.13 we get $|{\operatorname{SO}}_{2n}^{+}(q)|>(q^{2}-1)^{2}(q^{n}-1)q^{2n^{2}-2n-4}$. Multiplying out we see that $(q^{2}-1)^{2}(q^{n}-1)\geq(q-1)q^{n+3}$. Using that $q-1\geq q/2$ then achieves the proof. ∎
Lemma 4.22.
For $G={\operatorname{SO}}_{2n}^{\pm}(q)$ with $n\geq 4$ the subgroup $N$ has order $|N|\geq 2^{a-2}q^{n}$, where $n=ae+r$ with $0\leq r\leq e$ as introduced above.
Proof.
The relative Weyl group of a Sylow $d$-torus of $G$ is either $G(2e,1,a)$ or its normal subgroup $G(2e,2,a)$ of index 2. Now arguing as in the proof of Lemma 4.3 we see
$(q^{e}\pm 1)^{a}|G(2e,2,a)|\geq((3/4)q)^{e}(2e)^{a}a!/2\geq 2^{a-1}q^{ae}.$
Combining this with the bound in Lemma 4.21 for $|{\operatorname{SO}}_{2r}^{\pm}(q)|$ we conclude. ∎
Proposition 4.23.
Let $G=\Omega_{2n}^{\pm}(q)$ with $n\geq 4$, and assume $M$ is a maximal subgroup of $G$ in class $\mathcal{S}$} containing the normaliser of a Sylow $d$-torus. Then $M<G$ is one of:
  1. (1)
    ${\mathfrak{A}}_{9}<\Omega_{8}^{+}(2)$, with $d=3$, $\ell=7$;
  2. (2)
    $2.\Omega_{8}^{+}(2)<\Omega_{8}^{+}(3)$, with $d=4$, $\ell=5$;
  3. (3)
    ${\mathfrak{A}}_{16}<\Omega_{14}^{+}(2)$, with $d=3$, $\ell=7$; or
  4. (4)
    ${\mathfrak{A}}_{12}<\Omega_{10}^{-}(2)$, with $d=3$, $\ell=7$.
Proof.
Let $S=F^{*}(M)/{\mathbf{Z}}(F^{*}(M))$. As before, the case $n\leq 6$ can be discussed using the tables in [2, §8], which leads to Conclusions (1), (2) and (4), so we may assume $n\geq 7$. The sporadic groups are dealt with as in the proofs of Propositions 4.4 and 4.15 using [11, Tab.]. Next assume $S={\mathfrak{A}}_{m}$ is alternating. As earlier this implies $m\leq 2n+2$. Arguing exactly as in Proposition 4.15 we see that necessarily $q=2$ for both the plus- and the minus-type. Then again comparing the rank of a maximal elementary abelian $\ell$-subgroup $E$ of $S$ and the size of a maximal abelian subgroup containing it with the corresponding data in $G$, we find that $\ell\leq 11$ and $n\leq 13$. Going through the cases, and using that
[…]
and
${\mathfrak{A}}_{2n+1}\leq\begin{cases}\Omega_{2n}^{+}(2)&\text{if $n\equiv 0,3\pmod{4}$},\\ \Omega_{2n}^{-}(2)&\text{if $n\equiv 1,2\pmod{4}$}\end{cases}$
(see [12, p. 187]) we only arrive at the additional item (3) of the conclusion.
If $S$ is of Lie type in cross characteristic, we again use the lower bounds for faithful projective representations from [24] and argue as before. No further cases arise. The same holds for groups of Lie type in the same characteristic. ∎
Remark 4.24.
Observe that the maximal subgroups of type $\Omega_{8}^{+}(2)$ of $\Omega_{8}^{+}(3)$ are not invariant under the outer diagonal automorphism induced by ${\operatorname{SO}}_{8}^{+}(3)$ (see [2, Tbl. 8.50]) while a $4$-torus is. Since all semisimple classes are invariant under diagonal automorphisms, this means that $\Omega_{8}^{+}(2)$ cannot occur as a subnormaliser of a 5-element of ${\operatorname{SO}}_{8}^{+}(3)$.
We now determine the subnormalisers. The $d$-split Levi subgroups of ${\operatorname{GO}}_{2n}^{\epsilon}(q)$ have the form
${\operatorname{GL}}_{n_{1}}(q^{e})\times\cdots\times{\operatorname{GL}}_{n_{t}}(q^{e})\times{\operatorname{GO}}_{2s}^{\epsilon}(q)\qquad\text{if $d=e$ is odd},$
respectively
${\operatorname{GU}}_{n_{1}}(q^{e})\times\cdots\times{\operatorname{GU}}_{n_{t}}(q^{e})\times{\operatorname{GO}}_{2s}^{\epsilon^{\prime}}(q)\qquad\text{if $d=2e$ is even},\quad$
where $\epsilon^{\prime}=\epsilon(-1)^{\sum n_{i}}$ and $e\sum_{i}n_{i}+s=n$ in either case (see [9, Exmp. 3.5.15]) and we may assume $n_{1}\geq\cdots\geq n_{t}$. Here again we have $s\equiv r\pmod{e}$ and thus $s\geq r$. As before will write $L_{d}(n_{1},\ldots,n_{t};s)$ for such a $d$-split Levi subgroup.
Theorem 4.25.
Let $G={\operatorname{SO}}_{2n}^{+}(q)$ with $n\geq 4$, let $\ell\nmid q$ be a prime such that Sylow $\ell$-subgroups of $G$ are abelian and ${\mathbf{S}}_{d}\leq{\mathbf{G}}$ a Sylow $d$-torus where $d=e_{\ell}(q)$. Set $e:=e_{\ell}(q^{2})$ and let $a,r$ be as defined above. Then for $x\in{\mathbf{S}}_{d}^{F}$ an $\ell$-element we have one of
  1. (1)
    ${\operatorname{Sub}}_{G}(x)={\mathbf{N}}_{G}({\mathbf{S}}_{d})$ if ${\mathbf{C}}_{G}(x)={\mathbf{C}}_{G}({\mathbf{S}}_{d})$;
  2. (2)
    ${\operatorname{Sub}}_{G}(x)=\big({\operatorname{GO}}_{2ae}^{+}(q)\times{\operatorname{GO}}_{2r}^{+}(q)\big)\cap G$ if $r>0$, $2ae/d$ is even, ${\mathbf{C}}_{G}(x)=L_{d}(n_{1},\ldots,n_{t};r)$ with $n_{1}>1$;
  3. (3)
    ${\operatorname{Sub}}_{G}(x)=\big({\operatorname{GO}}_{2ae}^{-}(q)\times{\operatorname{GO}}_{2r}^{-}(q)\big)\cap G$ if $r>0$, $2ae/d$ is odd, ${\mathbf{C}}_{G}(x)=L_{d}(n_{1},\ldots,n_{t};r)$ with $n_{1}>1$;
  4. (4)
    $M=({\operatorname{GO}}_{2e}^{+}(q)\wr{\mathfrak{S}}_{a})\cap G$ if $r=0$, $a>1$, $d$ is odd and ${\mathbf{C}}_{G}(x)=L_{d}(1,\ldots,1;e)$;
  5. (5)
    $M=({\operatorname{GO}}_{2e}^{-}(q)\wr{\mathfrak{S}}_{a})\cap G$ if $r=0$, $a$ is even, $d$ is even and ${\mathbf{C}}_{G}(x)=L_{d}(1,\ldots,1;e)$;
  6. (6)
    $M={\operatorname{GL}}_{n}(q).2$ if $e=d=n/2$ is odd and ${\mathbf{C}}_{G}(x)=L_{d}(2;0)$;
  7. (7)
    $M={\operatorname{GU}}_{n}(q).2$ if $e=d/2=n/2$ is odd and ${\mathbf{C}}_{G}(x)=L_{d}(2;0)$;
  8. (8)
    $M={\operatorname{GO}}_{n}^{+}(q^{2}).2\cap G$ if $n=d\equiv 0\pmod{4}$ and ${\mathbf{C}}_{G}(x)=L_{d}(2;0)$; or
  9. (9)
    ${\operatorname{Sub}}_{G}(x)=G$ otherwise.
Proof.
With Proposition 2.2 we may assume Sylow $\ell$-subgroups are non-cyclic and so $a>1$. The only groups in Proposition 4.23 with non-cyclic Sylow $\ell$-subgroups are $2.\Omega_{8}^{+}(2)$ in $\Omega_{8}^{+}(3)$ with $\ell=5$, and ${\mathfrak{A}}_{16}$ in $\Omega_{14}^{+}(2)$ with $\ell=7$, but explicit computation shows that in either case the subnormalisers are as in (5) or (2) of the conclusion.
By [5, Prop. 2.2] the centraliser $C:={\mathbf{C}}_{G}(x)$ is $d$-split and thus $C=L_{d}(n_{1},\ldots,n_{t};s)$ as introduced above. We may assume $C$ is contained in one of the maximal subgroups $M$ of $G$ listed in Proposition 4.19 (as otherwise we have ${\operatorname{Sub}}_{G}(x)=G$ by Proposition 2.1). Also, if $C={\mathbf{C}}_{G}({\mathbf{S}}_{d})$ then we are in Case (1), so now suppose ${\mathbf{C}}_{G}({\mathbf{S}}_{d})$ is strictly contained in $C$. Assume first $r=0$. Then Cases (4)–(9) are relevant; note that in Cases (10) and (11) we have $\ell=3$ and Sylow $3$-subgroups of $G$ are non-abelian. Also, in Case (8) Sylow $\ell$-subgroups are cyclic, and similarly in Cases (6) and (7) unless $n=2e$. If $C$ lies in ${\operatorname{GO}}_{2e}^{\pm}(q)\wr{\mathfrak{S}}_{a}$ then necessarily $s=e$ and all $n_{i}=1$, as in (4) and (5) of the conclusion. If $C$ lies in ${\operatorname{GL}}_{n}(\pm q).2$ then clearly $s=0$ and the assumptions in Proposition 4.19(4) and (5) force $t=1$, $n_{1}=2$, as in (6) and (7) of the conclusion. If $C$ lies in ${\operatorname{GO}}_{n}^{+}(q^{2}).2$ then again we must have $s=0$ and $C=L_{d}(2;0)$ as in (8). Note that the conditions in (6), (7) and (8) are mutually exclusive.
Now assume $r>0$. If $C$ lies in ${\operatorname{GO}}_{2ae}^{\pm}(q)\times{\operatorname{GO}}_{2r}^{\pm}(q)$ then we must have $s=r$ and thus $n_{1}>1$ since $C>{\mathbf{C}}_{G}(x)$, as in (1) or (2) of our conclusion. If $q=2$, $r=1$ and $C$ lies in ${\operatorname{Sp}}_{2n-2}(2)$ then in fact we see $C$ lies in a subgroup ${\operatorname{GO}}_{2n-2}^{+}(2)=({\operatorname{GO}}_{2n-2}^{+}(2){\operatorname{GO}}_{2}^{+}(2))\cap G$ (see Proposition 4.11(7)) and we are back in (1) or (2). Finally, in all of the cases discussed above apart from the last one, ${\operatorname{Sub}}_{G}(x)$ must equal the relevant maximal subgroup, as can be seen by using our earlier descriptions of maximal subgroups of classical groups containing the normaliser of a Sylow $d$-torus in Propositions 4.1, 4.7, 4.11, 4.12, 4.19 and 4.20. ∎
Remark 4.26.
Observe that for $n=4$, cases (5) and (8) of Theorem 4.25 with $d=4$ are conjugate under triality (if $q$ is even). (See [2, Tab. 8.50].) Since a Sylow $d$-torus ${\mathbf{S}}_{d}$ of ${\mathbf{G}}={\operatorname{SO}}_{8}$ can be chosen invariant under a triality automorphism $\tau$ commuting with $F$, this shows that subnormalisers can behave strangely with respect to upward extensions, even by cyclic groups: if $x\in{\mathbf{S}}_{d}^{F}$ has centraliser $L_{4}(2;0)$, for example, then its subnormaliser in $G={\operatorname{SO}}_{8}^{+}(q)$ is as given in (6), (7) or (8) of Theorem 4.25, but its subnormaliser in $\hat{G}:=G\langle\tau\rangle$ is $\hat{G}$. A similar phenomenon for 3-elements in $^{2}\!F_{4}(2)^{\prime}<{}^{2}\!F_{4}(2)$ was already observed in [16, §5.3].
We complete our investigations by considering subnormalisers in orthogonal groups of minus-type.
Theorem 4.27.
Let $G={\operatorname{SO}}_{2n}^{-}(q)$ with $n\geq 4$, let $\ell\nmid q$ be a prime such that Sylow $\ell$-subgroups of $G$ are abelian and ${\mathbf{S}}_{d}\leq{\mathbf{G}}$ a Sylow $d$-torus where $d=e_{\ell}(q)$. Set $e:=e_{\ell}(q^{2})$ and let $a,r$ be as defined above. Then for $x\in{\mathbf{S}}_{d}^{F}$ an $\ell$-element we have one of
  1. (1)
    ${\operatorname{Sub}}_{G}(x)={\mathbf{N}}_{G}({\mathbf{S}}_{d})$ if ${\mathbf{C}}_{G}(x)={\mathbf{C}}_{G}({\mathbf{S}}_{d})$;
  2. (2)
    ${\operatorname{Sub}}_{G}(x)=\big({\operatorname{GO}}_{2ae}^{+}(q)\times{\operatorname{GO}}_{2r}^{-}(q)\big)\cap G$ if $2ae/d$ is even and ${\mathbf{C}}_{G}(x)=L_{d}(n_{1},\ldots,n_{t};r)$ with $n_{1}>1$;
  3. (3)
    ${\operatorname{Sub}}_{G}(x)=\big({\operatorname{GO}}_{2ae}^{-}(q)\times{\operatorname{GO}}_{2r}^{+}(q)\big)\cap G$ if $2ae/d$ is odd and ${\mathbf{C}}_{G}(x)=L_{d}(n_{1},\ldots,n_{t};r)$ with $n_{1}>1$;
  4. (4)
    $M=({\operatorname{GO}}_{2e}^{-}(q)\wr{\mathfrak{S}}_{a})\cap G$ if $r=0$, $a$ is odd, $d$ is even and ${\mathbf{C}}_{G}(x)=L_{d}(1,\ldots,1;e)$; or
  5. (5)
    ${\operatorname{Sub}}_{G}(x)=G$ otherwise.
Proof.
The argument is similar to but easier than the one for the orthogonal groups of plus-type. With the usual reductions we may assume $a\geq 2$ and $C:={\mathbf{C}}_{G}(x)$ is a $d$-split Levi subgroup of type $L_{d}(n_{1},\ldots,n_{t};s)$ lying in one of the maximal subgroups $M$ in (1)–(4) of Proposition 4.20. If $r=0$, so we are in Case (4), the containment $C\leq M$ forces $s=e$ and $n_{1}=\ldots=n_{t}=1$, so we reach (4) of the conclusion. If $r>0$ then $C\leq M$ gives $s=r$ and we are in (2) or (3) of the conclusion; note that again Case (3) of Proposition 4.20 does not appear as a subnormaliser since in that case $C$ is contained in a proper subgroup ${\operatorname{GO}}_{2n-2}^{-}(2)=({\operatorname{GO}}_{2n-2}^{-}(2){\operatorname{GO}}_{2}^{+}(2))\cap G$ of ${\operatorname{Sp}}_{2n-2}(2)$ (see Proposition 4.11(8)). ∎
Looking back on the results obtained for the various types of classical groups we observe that, as in the case of exceptional groups, the subnormalisers of semisimple $\ell$-elements (in abelian Sylow subgroups) occurring in classical groups are always normalisers of suitable subsystem subgroups, with the sole exception of the extension field subgroups of ${\operatorname{SO}}_{2n}^{+}(q)$ in Theorem 4.25(8), very similar to the situation for semisimple elements in algebraic groups in [16, Thm 6.8]. We do not see, though, how to deduce this a priori.
5. Subnormalisers in symmetric groups
In this section we describe subnormalisers of $p$-elements in symmetric groups ${\mathfrak{S}}_{n}$ with abelian Sylow $p$-subgroups. The results are very similar to those for the special linear groups but easier to show. We write $n=ap+r$ with $0\leq r<p$, where have $a\leq p-1$ as otherwise the Sylow $p$-subgroups of ${\mathfrak{S}}_{n}$ are non-abelian. Note that a Sylow $p$-subgroup of ${\mathfrak{S}}_{n}$ is then elementary abelian, and the conjugacy class of a $p$-element in ${\mathfrak{S}}_{n}$ is uniquely determined by its number of cycles of length $p$.
Proposition 5.1.
Let $G={\mathfrak{S}}_{n}$, $p$ a prime and $n=ap+r$ with $a,r\leq p-1$. Let $x\in G$ have cycle type $(p)^{k}$. Then
${\operatorname{Sub}}_{G}(x)=\begin{cases}\big(C_{p}.C_{p-1}\big)\wr{\mathfrak{S}}_{a}\times{\mathfrak{S}}_{r}&\text{if $k=a$ ($x$ is picky)},\\ {\mathfrak{S}}_{p}\wr{\mathfrak{S}}_{a}&\text{if $r=0$ and $k=a-1\geq 1$},\\ G&\text{otherwise}.\end{cases}$
If $p>2$ and so $x\in{\mathfrak{A}}_{n}$ then ${\operatorname{Sub}}_{{\mathfrak{A}}_{n}}(x)={\operatorname{Sub}}_{G}(x)\cap{\mathfrak{A}}_{n}$.
Proof.
A Sylow $p$-subgroup $P$ of ${\mathfrak{S}}_{n}$ has normaliser $\big(C_{p}.C_{p-1}\big)\wr{\mathfrak{S}}_{a}\times{\mathfrak{S}}_{r}$, while the centraliser of $x$ has the form $C_{p}\wr{\mathfrak{S}}_{k}\times{\mathfrak{S}}_{n-kp}$. Thus, if $x\in P$ and $k=a$ then ${\mathbf{C}}_{G}(x)\leq{\mathbf{N}}_{G}(P)$ and hence $x$ is picky by [16, Cor. 2.7 and Prop. 2.12]. Now assume $k<a$. If $r=0$ and k=a1k=a-1 then clearly, ${\mathbf{C}}_{G}(x)$ and ${\mathbf{N}}_{G}(P)$ are both contained in a subgroup $M={\mathfrak{S}}_{p}\wr{\mathfrak{S}}_{a}$, and in fact, $M$ is generated by these, so we are in case (2) of the conclusion. So finally assume k<a1k<a-1, or $r>0$ and k=a1k=a-1. Then ${\mathbf{C}}_{G}(x)$ contains a symmetric group ${\mathfrak{S}}_{2p}$, respectively ${\mathfrak{S}}_{r+p}$, and ${\operatorname{Sub}}_{G}(x)=\langle{\mathbf{C}}_{G}(x),{\mathbf{N}}_{G}(P)\rangle$ acts transitively on $\{1,\ldots,n\}$, which forces ${\operatorname{Sub}}_{G}(x)=G$ by Jordan’s theorem.
Now assume $p>2$. Clearly, ${\operatorname{Sub}}_{{\mathfrak{A}}_{n}}(x)\leq{\operatorname{Sub}}_{G}(x)\cap{\mathfrak{A}}_{n}$. For the converse, we may assume $p>3$ by explicit computation, and then the same reasoning as for ${\mathfrak{S}}_{n}$ applies. ∎
Note that for $n=6$, $p=3$, the first two cases in Proposition 5.1 are conjugate under the exceptional outer automorphism.
The fact that $x$ is picky for $k=a$ was already shown by Maróti, Martínez Madrid and Moretó [19, Thm 3.9]. The situation for primes $p\leq\sqrt{n}$ is much more involved and many different types of subnormalisers can arise. For the prime 2 they were completely determined by Martínez Madrid [20]. For example in ${\mathfrak{S}}_{15}$ there are eight different classes of subnormalisers of 2-elements.
6. On subnormalisers in sporadic groups
The Sylow $p$-subgroups in sporadic simple groups $G$ are cyclic of prime order in most cases, and then the subnormaliser of any non-trivial $p$-element $x\in G$ is just ${\mathbf{N}}_{G}(\langle x\rangle)$, by [16, Prop. 2.12]. Here we consider the remaining instances of abelian Sylow subgroups:
Proposition 6.1.
Let $G$ be a sporadic simple group and $p$ a prime such that Sylow $p$-subgroups of $G$ are abelian, but not cyclic of prime order. Let $x\in G$ be a $p$-element. Then ${\operatorname{Sub}}_{G}(x)=G$ unless $(G,p)$ are as given in Table 4.
Table 4. Subnormalisers in sporadic groups
[…]
Proof.
The sporadic groups $G$ satisfying the assumptions of the proposition are $J_{1}$ for $p=2$, $M_{11},M_{22},M_{23},HS$ for $p=3$, $J_{2},Suz,He,Fi_{22},Fi_{23},Fi_{24}^{\prime}$ for $p=5$, $Co_{1},Th,B$ for $p=7$, and the monster group for $p=11$. For the smaller groups the claim is easily verified by a computer calculation. We just comment on the larger cases (variations of the given arguments would in fact also allow to treat the smaller cases by hand). For $G=Fi_{24}^{\prime}$ with $p=5$ there is a unique class of 5-elements. Now a Sylow 5-subgroup of $G$ is contained in a maximal subgroup $M=Fi_{23}$, where the subnormaliser of any 5-element $x$ equals $M$. Since $|{\mathbf{C}}_{G}(x)|>|{\mathbf{C}}_{M}(x)|$ this shows ${\operatorname{Sub}}_{G}(x)=G$ by Proposition 2.1. For $G=Th$ with $p=7$ there is again a unique class of 7-elements. Here a Sylow 7-subgroup $P$ of $G$ is contained in a maximal subgroup $M={}^{3}\!D_{4}(2).3$, and this group has one class of 7-elements $x$ with ${\operatorname{Sub}}_{M}(x)=M$ (see Theorem 3.2). Since $|{\mathbf{N}}_{G}(P)|>|{\mathbf{N}}_{M}(x)|$ we conclude ${\operatorname{Sub}}_{G}(x)=G$ (using again Proposition 2.1). For $G=B$ with $p=7$ we use our previous result on the maximal subgroup $M=Th$ by noting that centralisers of 7-elements in $G$ are larger than in $M$. Finally assume $G$ is the monster group and $p=11$. Here $G$ has a unique class of 11-elements, but the centraliser of such an element has order not dividing the order of a Sylow 11-normaliser, which is maximal in $G$, so again ${\operatorname{Sub}}_{G}(x)=G$.
Since $G=Fi_{22}$ is a maximal subgroup of $^{2}\!E_{6}(2)$, the subnormalisers of the unique class of 5-elements $x\in G$ can be derived from Theorem 3.2 for an upper bound and Remark 4.26 for a lower bound. ∎
The subnormaliser $J_{2}.2$ in $Suz$ of course contains the subnormaliser $J_{2}$ in the maximal subgroup $G_{2}(4)$ of $Suz$ found in Theorem 3.2.
Proof of Theorem 2.
Let $S$ be quasi-simple. By [16, Lemma 2.15] we may assume ${\mathbf{Z}}(S)=1$, so $S$ is simple. For $S$ a sporadic group, the subnormalisers were found in Proposition 6.1, while for $S={\mathfrak{A}}_{n}$ they are known by Proposition 5.1, respectively [20] when $p=2$.
So finally assume $S$ is simple of Lie type. If $p$ is the defining characteristic of $S$, then $S\cong{\operatorname{L}}_{2}(p^{f})$ and the non-trivial $p$-elements are picky by [19, Thm 6.1] or [16, Prop. 3.6]. If $p$ is not the defining characteristic, then either $p>2$ or again $S={\operatorname{L}}_{2}(q)$. In the former case, subnormalisers are determined in Sections 3 and 4. So let $S={\operatorname{L}}_{2}(q)$ with $q\equiv 3,5\pmod{8}$. The cases when a 2-element $x\in S$ is picky are described in [19, Thm 6.1] and [17, Lemma 3.7]. Otherwise, ${\operatorname{Sub}}_{S}(x)$ properly contains the normalisers ${\mathfrak{A}}_{4}$ of a Sylow 2-subgroup and of a maximal torus of order (q1)/2(q-1)/2 or $(q+1)/2$ and hence equals $S$. ∎